### v2

# A Basic Evaluation of Neural Networks Trained with the Error Diffusion Learning Algorithm

Abstract
Purpose
This paper presents a contemporary formulation of Kaneko’s Error Diffusion Learning Algorithm (EDLA) framework and evaluates its effectiveness through parity check, regression, and image classification tasks. The EDLA is a biologically inspired learning algorithm that is an alternative to the conventional backpropagation algorithm for training artificial neural networks.
Design/methodology/approach
EDLA employs a single global error signal diffused across neural networks comprising paired excitatory-inhibitory sublayers, eliminating traditional layer-wise error backpropagation. This study evaluates EDLA’s effectiveness using benchmark tasks—parity check, regression, and image classification—by systematically varying the neuron count, network depth, and learning rates to assess its performance comprehensively.
Findings
The experimental results demonstrate that EDLA achieves consistently high accuracy across multiple benchmarks, highlighting its effectiveness as a training algorithm for neural networks. The choice of learning rate, neuron count, and network depth significantly influences EDLA’s efficiency and convergence speed. Analysis of internal network representations reveals meaningful feature extraction capabilities comparable to networks trained via traditional backpropagation.
Originality/value
This study introduces EDLA, a biologically plausible alternative to traditional backpropagation previously underrecognized due to language barriers. By providing a modern and comprehensive formulation of EDLA, systematically evaluating its performance, and presenting empirical evidence of its effectiveness, this study significantly enhances the visibility and accessibility of EDLA, contributing substantially to biologically inspired neural network training methodologies.
1 Introduction
Artificial neural networks have substantially advanced numerous research domains, including image processing, signal processing, and natural language processing. The remarkable success of these networks largely stems from the backpropagation method proposed by Rumelhart et al. (Rumelhart et al., 1986), which has become a standard method for training neural networks across diverse architectures.
The theoretical foundation of neural networks can be traced back to the linear threshold unit (neuron) introduced by McCulloch and Pitts (Kleene, 1956; Mcculloch and Pitts, 1943). The fundamental mechanism of this neuron is based on the all-or-none law (Adrian, 1914; Kato, 1926) and still underlies modern neural network architectures. Rosenblatt’s development of the perceptron (Rosenblatt, 1958) represents another critical milestone in neural networks, marking the first realization of a trainable artificial neuron. However, a single-layer structure of perceptrons does not inherently have the capability to create nonlinear decision boundaries. Although multilayer perceptrons (MLPs) resolve this limitation, training these deeper networks has been a substantial challenge. The introduction of backpropagation has provided a solution for training MLPs. Subsequently, it has become the dominant paradigm for training various deep neural network architectures.
Despite the success of backpropagation, its biological plausibility has been questioned (Crick, 1989; Mostafa et al., 2018). This gap between artificial and biological neural systems has motivated the development of alternative learning algorithms. One example is Kaneko’s Error Diffusion Learning Algorithm (EDLA) (Kaneko, 1999). EDLA introduces a biologically inspired architecture that incorporates both excitatory and inhibitory synapses, reflecting biological neural systems more closely. The distinctive feature of this algorithm is its error-diffusion mechanism. In this approach, the difference between the network output and target values is diffused throughout the network during training, resulting in efficient training of sigmoid neural networks. Unlike standard backpropagation, the EDLA diffuses only a single error signal throughout the network, eliminating the need for backward propagation and error calculation for each layer. This simplification addresses one of the key problems highlighted by Mostafa (2018), which artificial neural networks based on backpropagation require buffering past presynaptic activities and synchronizing them with the corresponding error signals propagated sequentially from higher layers.
This study provides a comprehensive overview of the EDLA framework and evaluates its performance on several benchmark tasks including parity checks, regression, and image classification. Through these evaluations, we aim to promote the EDLA framework and facilitate the development of training methods for biologically plausible neural networks.
Portions of this text were previously published as part of a preprint (https://arxiv.org/abs/2504.14814).
2 Error Diffusion Learning Algorithm (EDLA)
The Error Diffusion Learning Algorithm (EDLA) is a novel approach to neural network learning initially proposed by Kaneko (Kaneko, 1999). Although the original model is available only in Japanese, this section provides a comprehensive reinterpretation of EDLA.
2.1 Network Architecture
Figure 1 illustrates the neural network architecture employed by the EDLA. Like traditional MLPs, EDLA adopts a multilayer structure comprising neurons and synapses. However, EDLA incorporates two types of neurons: positive neurons (p-neurons) and negative neurons (n-neurons). The synaptic connections are categorized based on the types of neurons they link: excitatory synapses connect neurons of identical type, whereas inhibitory synapses connect neurons of different types. A key characteristic of the EDLA is that it has only one output neuron.
Figure 1: Schematic representation of the Error Diffusion Learning Algorithm (EDLA) network architecture. The network consists of multiple layers of positive neurons (white circles) and negative neurons (gray circles). The red and blue small circles at the connections indicate excitatory and inhibitory synapses, respectively.
2.2 Learning Rule and Mathematical Formulation
The input to the EDLA network is represented by the vector $\mathbf{x}=[x_{0},x_{1},\ldots,x_{N}]^{T}\in\mathbb{R}^{N+1}$, where $x_{0}$ corresponds to the bias set to one, and $N$ is the number of input features. In the EDLA architecture, each layer consists of positive and negative sublayers. The outputs from the positive and negative sublayers of the input layer are denoted by $\mathbf{x}^{p}$ and $\mathbf{x}^{n}$, respectively. For the input layer, $\mathbf{x}^{p}=\mathbf{x}^{n}=\mathbf{x}$.
For a neuron $j$ in the first hidden layer $(l=1)$, the activations for the positive and negative neurons, $a_{j}^{p,(1)}$ and $a_{j}^{n,(1)}$, are computed as follows:
$a_{j}^{p,(1)}={\mathbf{w}_{j}^{pp,(1)}}^{T}\mathbf{x}^{p}+{\mathbf{w}_{j}^{pn,(1)}}^{T}\mathbf{x}^{n},$ (1)
$a_{j}^{n,(1)}={\mathbf{w}_{j}^{np,(1)}}^{T}\mathbf{x}^{p}+{\mathbf{w}_{j}^{nn,(1)}}^{T}\mathbf{x}^{n}.$ (2)
In these equations, $\mathbf{w}_{j}^{pp,(1)}=[w_{j0}^{pp},\ldots,w_{jN}^{pp}]^{T}$ represents the vector of synaptic weights connecting the positive neurons in the input layer to positive neuron $j$ in the first hidden layer $(l=1)$. Similarly, $\mathbf{w}_{j}^{pn,(1)}$, $\mathbf{w}_{j}^{np,(1)}$, and $\mathbf{w}_{j}^{nn,(1)}$ denote the weight vectors for the other combinations of neuron types in the connections between the input and the first hidden layers.
The outputs of the positive and negative neurons $j$ in the first hidden layer are given by:
$z_{j}^{p,(1)}=g(a_{j}^{p,(1)})\quad\text{and}\quad z_{j}^{n,(1)}=g(a_{j}^{n,(1)})$ (3)
where $g(\cdot)$ is the activation function. In subsequent hidden layers $(l=2,...,L)$ where $L$ refers to the final hidden layer, the neuron outputs are given by
$z_{j}^{p,(l)}=g({\mathbf{w}_{j}^{pp,(l)}}^{T}\mathbf{z}^{p,(l-1)}+{\mathbf{w}_{j}^{pn,(l)}}^{T}\mathbf{z}^{n,(l-1)}),$ (4)
$z_{j}^{n,(l)}=g({\mathbf{w}_{j}^{np,(l)}}^{T}\mathbf{z}^{p,(l-1)}+{\mathbf{w}_{j}^{nn,(l)}}^{T}\mathbf{z}^{n,(l-1)}).$ (5)
The final output of the EDLA network $y$ is computed as
$y=g({\mathbf{w}^{pp,(L+1)}}^{T}\mathbf{z}^{p,(L)}+{\mathbf{w}^{pn,(L+1)}}^{T}\mathbf{z}^{n,(L)}).$ (6)
Synaptic weights within the EDLA network are constrained to maintain specific relationships, reflecting the excitatory and inhibitory nature of the connections:
$w^{pp,(l)}_{ji}\geq 0,\quad w^{pn,(l)}_{ji}\leq 0,\quad w^{np,(l)}_{ji}\leq 0,\quad w^{nn,(l)}_{ji}\geq 0.$ (7)
The objective of training an EDLA network is to identify the optimal parameter set $\mathbf{W}$ that minimizes a predefined error function $E(\mathbf{W})$, similar to traditional standard MLPs. Gradient descent is a widely used optimization method in which parameters are iteratively updated in a direction that locally reduces the value of $E(\mathbf{W})$. In practice, most practitioners use stochastic gradient descent (SGD) (LeCun et al., 2015), which computes gradient estimates efficiently using subsets of training data (mini-batches). The EDLA also employs a method based on SGD.
Although the EDLA utilizes the principles of gradient descent, it incorporates an approximation in computing the errors used for weight updates. In an MLP trained by backpropagation, the error $\delta_{j}^{(l)}$ for neuron $j$ in hidden layer $l$ is computed as:
$\delta_{j}^{(l)}=g^{\prime}(a_{j}^{(l)})\sum_{k}w_{jk}^{(l+1)}\delta_{k}^{(l+1)},$ (8)
where $g^{\prime}(a_{j}^{(l)})$ is the derivative of the activation function. In contrast, the EDLA approximates these errors using a single global error signal $d$, defined as
$d=-\frac{\partial E(\mathbf{W})}{\partial y}.$ (9)
This global error is uniformly diffused throughout the network. Specifically, the approximated errors, $\tilde{\delta}^{pp}_{ji}$, $\tilde{\delta}^{np}_{ji}$, $\tilde{\delta}^{nn}_{ji}$, and $\tilde{\delta}^{pn}_{ji}$ in the EDLA, are expressed as:
$\tilde{\delta}^{pp,(l)}_{ji}=dg^{\prime}(a^{p,(l)}_{j}),\tilde{\delta}^{np,(l)}_{ji}=dg^{\prime}(a^{n,(l)}_{j}),$ (10)
$\tilde{\delta}^{nn,(l)}_{ji}=dg^{\prime}(a^{n,(l)}_{j}),\tilde{\delta}^{pn,(l)}_{ji}=dg^{\prime}(a^{p,(l)}_{j}).$ (11)
A key distinction between EDLA and traditional backpropagation is EDLA’s simplified approximation of error computations. Whereas traditional backpropagation individually computes and propagates neuron-specific errors $\delta$s from subsequent layers, EDLA diffuses only one global error signal $d$ uniformly across all layers. This diffusion mechanism significantly reduces complexity by eliminating the need for individual neuron-level error calculations, highlighting a unique feature of the EDLA approach compared with traditional backpropagation methods.
The weight updates in EDLA differ depending on the sign of the global error $d$. When $d>0$, the weights originating from the positive neurons $w^{pp,(l)}_{ji}$ and $w^{np,(l)}_{ji}$ are updated as follows:
$w^{pp,(l)}_{ji}\leftarrow w^{pp,(l)}_{ji}+\eta\tilde{\delta}^{pp,(l)}_{ji}z^{p,(l-1)}_{i}=w^{pp,(l)}_{ji}+\eta dg^{\prime}(a^{p,(l)}_{j})z^{p,(l-1)}_{i},$ (12)
$w^{np,(l)}_{ji}\leftarrow w^{np,(l)}_{ji}-\eta\tilde{\delta}^{np,(l)}_{ji}z^{p,(l-1)}_{i}=w^{np,(l)}_{ji}-\eta dg^{\prime}(a^{n,(l)}_{j})z^{p,(l-1)}_{i}.$ (13)
Conversely, when $d<0$, the weights originating from the negative neurons, $w^{nn,(l)}_{ji}$ and $w^{pn,(l)}_{ji}$ are updated as:
$w^{nn,(l)}_{ji}\leftarrow w^{nn,(l)}_{ji}-\eta\tilde{\delta}^{nn,(l)}_{ji}z^{n,(l-1)}_{i}=w^{nn,(l)}_{ji}-\eta dg^{\prime}(a^{n,(l)}_{j})z^{n,(l-1)}_{i},$ (14)
$w^{pn,(l)}_{ji}\leftarrow w^{pn,(l)}_{ji}+\eta\tilde{\delta}^{pn,(l)}_{ji}z^{n,(l-1)}_{i}=w^{pn,(l)}_{ji}+\eta dg^{\prime}(a^{p,(l)}_{j})z^{n,(l-1)}_{i}.$ (15)
where $\eta$ denotes the learning rate. Thus, the global error signal directly determines whether the positive or negative neuron weights are adjusted during the training.
The behavior of the EDLA becomes intuitively clear when employing the sigmoid activation function $g(a)=1/(1+\exp(-a))$ along with the squared error loss function defined as $E(\mathbf{w})=\frac{1}{2}(y-t)^{2}$. The global error signal $d$ is represented as follows:
$d=-\frac{\partial E(\mathbf{w})}{\partial y}=-\frac{\partial}{\partial y}\left[\frac{1}{2}(y-t)^{2}\right]=t-y.$ (16)
When the network output is smaller than the target value ($d>0$), the absolute values of the weights originating from positive neurons increase due to the properties $g^{\prime}(a)>0$ and $z>0$. Such weight adjustments amplify the outputs of positive neurons in subsequent layers while simultaneously suppressing the outputs of negative neurons in those layers. Consequently, these modifications reduce the discrepancy between the output and target value of the network. Conversely, when the network output exceeds the target value ($d<0$), the absolute values of the weights originating from negative neurons increase. These weight adjustments decrease the outputs of positive neurons in subsequent layers while simultaneously increasing the outputs of negative neurons. Thus, regardless of the sign of the global error signal, EDLA systematically reduces the difference between the network output and target values through successive weight updates. Importantly, this characteristic learning behavior of EDLA is maintained as long as $g^{\prime}(a)>0$, $z>0$, and the sign of the derivative of the loss function properly encodes whether the network output is below or above the target value.
For the EDLA, the activation function is typically chosen as the sigmoid function. It constrains the output of the neuron between 0 and 1 and its derivative between 0 and 0.25. By limiting both the neuron outputs and their derivatives, the sigmoid activation function provides stability to the training dynamics, avoiding huge weight updates and facilitating stable convergence.
2.3 Biological Interpretation
In the EDLA framework, the global error signal $d$ functions as a plasticity-guiding neuromodulatory signal, orchestrating synaptic weight updates through the coordinated activities of the presynaptic and postsynaptic neurons. This concept is strongly reminiscent of Hebbian learning (Hebb, 1949), which posits that changes in synaptic strength are driven by correlated patterns of neuronal activation. Specifically, the EDLA learning rule, represented by $\eta dg^{\prime}(a^{p,(l)}_{j})z^{p,(l-1)}_{i}$ for synaptic weights $w^{pp,(l)}_{ji}$, is proportional to the product of the derivative of the postsynaptic neuron’s activation function $g^{\prime}(a^{p,(l)}_{j})$ and the output of the presynaptic neuron $z^{p,(l-1)}_{i}$. This mathematical relationship can be interpreted as a Hebbian-like plasticity mechanism, wherein weight adjustments reflect a direct interaction between a presynaptic neuron (characterized by $z^{p,(l-1)}_{i}$ and a postsynaptic neuron (captured by $g^{\prime}(a^{p,(l)}_{j})$).
Moreover, EDLA employs a global error signal diffusion mechanism, in which the error signal $d$ uniformly propagates throughout the entire network. This diffusion process is conceptually analogous to biological neuromodulation and closely parallels the reward prediction error (RPE) mechanism observed in biological neural systems, which guides synaptic plasticity across neural circuits through diffuse neuromodulatory signals (Roelfsema and Holtmaat, 2018). This neuromodulatory-inspired global diffusion effectively orchestrates all synaptic modifications in the network; thereby, this mirrors biological processes wherein neuromodulators such as dopamine broadly regulate neural plasticity based on RPE signals.
2.4 Limitation
The inherent simplicity and approximation underlying EDLA introduce significant constraints, particularly concerning the network’s capability to handle multiple outputs effectively. In the EDLA framework, the global error signal $d$ is uniformly diffused throughout the network architecture. This global error diffusion inherently directs the optimization process toward minimizing the error associated with a single output node. Consequently, when dealing with single-output scenarios, the EDLA can efficiently and accurately specialize the network weights for a specific output. However, when an EDLA network contains multiple outputs, each output generates its own global error signal, which propagates throughout the network. Such simultaneous diffusion of distinct error signals leads to conflicting weight adjustments, as the network simultaneously attempts to minimize errors for multiple outputs. These conflicting updates inhibit the network’s ability to extract and generalize meaningful features from the input data, thereby undermining learning efficacy.
2.5 Extending EDLA to Multiple Outputs
The inherent single-output constraint of the EDLA represents a significant limitation when addressing tasks that require multiple outputs. A simple approach to adapting EDLA to overcome this limitation involves constructing multiple EDLA networks arranged in parallel. This parallelized architecture addresses a task that requires $K$ outputs by employing $K$-independent EDLA networks. Each individual network retains the fundamental EDLA architecture and operates autonomously. Input vector $\mathbf{x}$ is simultaneously provided to all networks, ensuring that they process the same input data. Each network independently computes output $y_{k}$, where $k\in\{1,\ldots,K\}$. In classification tasks, if the output $y_{k}$ is the maximum value among the outputs of the EDLA networks, $k$ is regarded as the overall output label.
The weights of each individual EDLA network are independently optimized using the standard EDLA update rule. Consequently, this parallel architecture retains the capability, simplicity, and core principles inherent to single-output EDLA models, effectively extending EDLA’s capability to handle multioutput tasks. However, this parallelization approach increases computational costs because the total number of EDLA networks scales linearly with the required number of outputs. This trade-off shows the challenge of maintaining the computational simplicity of single-output EDLA networks while contending with the additional resource demands necessary for multioutput learning.
Employing multiple single-output networks in an EDLA framework may seem computationally inefficient compared with traditional multi-output MLP architectures. Nevertheless, this parallelization method is reminiscent of module structures, which are organizational principles observed in biological neural systems, particularly in the brain (Casanova and Casanova, 2018). Mountcastle’s early neurophysiological research (Mountcastle, 1957) has demonstrated that neurons within the primate somatosensory cortex are organized into vertically aligned columns, each dedicated to a particular modality or receptive field. Additionally, neurons in the primary visual cortex are organized into modular structures known as hypercolumns, each specializing in processing specific features of visual stimuli such as orientation (Hubel and Wiesel, 1968). Each hypercolumn functions semi-independently, collectively contributing to overall visual perception by handling distinct components of the visual scene. Similarly, the inferior temporal cortex exhibits clusters of neurons that are selectively responsive to complex visual features including distinct object shapes (Sato et al., 2008; Tanaka, 1996). This clustered arrangement enables the efficient processing of diverse visual elements while simultaneously facilitating effective integration across features and categories. Another notable example is the barrel cortex of rodents, in which neural clusters (called barrels) are dedicated to independently processing tactile inputs from individual whiskers (Li and Crair, 2011; Woolsey and Van der Loos, 1970). Each barrel independently processes sensory information corresponding to its associated whisker, and the combined neural activity across the barrels generates a comprehensive tactile representation. These modular and distributed neural processing strategies enable the brain to handle complex and heterogeneous tasks efficiently while preserving both flexibility and robustness. Although the parallelized architecture of EDLA, which consists of multiple single-output networks dedicated to specific outputs, may be computationally less efficient than traditional multi-output MLPs, it reflects these biological modular architectures.
3 Experimental Setup
The neural networks are implemented in Python using NumPy for linear algebra operations and PyTorch for neural networks. The source code is available on GitHub at https://github.com/KazuhisaFujita/EDLA for reproducibility and further extension of this work.
4 Experimental Results
4.1 Parity Check Task
To evaluate the efficacy of the Error Diffusion Learning Algorithm (EDLA), we initially assessed its performance on a parity check task, a classical benchmark in neural network learning. This task involves determining whether the number of ones in a binary input vector is even or odd. The parity check task is a two-class classification problem. The EDLA network consists of input, hidden, and output layers. The neurons in the hidden layers use the sigmoid or ReLU activation functions (AFs). The output neuron is singular and employs the sigmoid AF, producing a value between 0 and 1. The output is interpreted as the probability of the input vector containing an odd number of ones. Thus, the output neuron is expected to yield a value above 0.5 for odd counts and below 0.5 for even counts. The total number of possible input combinations is $2^{n_{\mathrm{bit}}}$, where $n_{\mathrm{bit}}$ denotes the number of bits in the input vector. In this experiment, $n_{\mathrm{bit}}=2,3,4,5$ with each dataset including all possible combinations for a given $n_{\text{bit}}$. A mini-batch size of 4 is consistently employed across all configurations.
Figure 2 illustrates the performance of the EDLA network with a single hidden layer and the sigmoid AF across various learning rates (0.1, 1.0, 10.0) and neuron counts per layer (8, 16, 32, 64, 128). The results demonstrate that the EDLA network achieves high accuracy on the parity check task. The network attains perfect accuracy across all cases except for configurations with eight neurons when $n_{\mathrm{bit}}=5$. The number of epochs required to reach an accuracy of 0.9 decreases as the number of neurons increases. For learning rates of 10 and 1.0, the convergence time (reach epoch) is smaller than the learning rate of 0.1. Furthermore, for $n_{\mathrm{bit}}\in\{3,5\}$, the reach epoch at a learning rate of 1.0 is nearly identical to that at a learning rate of 10. For $n_{\mathrm{bit}}=2$, the reach-epoch for a learning rate of 1.0 is larger than that at a learning rate of 10, but for $n_{\mathrm{bit}}=4$ the reach epoch at a learning rate of 1.0 is smaller than that at a learning rate of 10. These results suggest that learning rates of 10 and 1.0 enable efficient training.
Figure 3 presents an analysis of the performance of the EDLA network using the sigmoid AF across various hidden layers (1, 2, 4, and 8 hidden layers) and neuron counts per layer with a fixed learning rate of 1.0. The results reveal that the accuracy consistently approaches or reaches 1.0 across different bit numbers ($n_{\mathrm{bit}}$), regardless of the number of hidden layers. However, the number of epochs required to achieve 0.9 accuracy reveals more nuanced dependencies on the network depth and neuron count. Specifically, for $n_{\mathrm{bit}}=2$, shallow networks (one layer) exhibit a relatively low reach epoch for smaller neuron numbers, but the reach epoch increased for larger neuron counts. The deeper networks do not exhibit significant reach epoch reductions for $n_{\mathrm{bit}}=2$. Conversely, for $n_{\mathrm{bit}}\in\{3,4\}$ with deeper architectures (four and eight layers), the reach epoch increases with the number of neurons. For $n_{\mathrm{bit}}=5$, the reach epoch decreases notably with the neuron count and stabilized for neuron counts of 64 and above. These results indicate that EDLA networks with sigmoid AF can consistently achieve high accuracy on the parity check task, with optimal performance depending on the complexity of the task and the chosen network configuration.
Figure 4 shows the performance of the EDLA network with a single hidden layer and the ReLU AF across various learning rates (0.01, 0.1, 1.0) and neuron counts of the hidden layer. For learning rates of 0.01 and 0.1, the network achieves high accuracy on the parity check task, with the number of epochs required to reach an accuracy of 0.9, decreasing as the number of neurons increases. For a learning rate of 0.01, the network exhibits a slower convergence speed than a learning rate of 0.1, particularly for smaller neuron counts. For a learning rate of 1.0, the performance is worse than that for the other learning rates. In particular, for a learning rate of 1.0 and $n_{\mathrm{bit}}\in\{4,5\}$, the network fails to achieve an accuracy of 0.9 even with a large number of neurons. This shows that a learning rate of 1.0, which is too high for the EDLA network with the ReLU, leads to instability in training and poor performance.
Figure 5 shows the accuracy and convergence speed of the EDLA network with the ReLU across various neuron counts and hidden layer depths. Both the accuracy and convergence speed improve with increasing neuron counts in the ReLU network. However, this improvement in the convergence speed does not continue indefinitely and eventually plateaus. These results indicate a performance comparable to that of sigmoid-based networks, with a consistently faster convergence observed for ReLU-based configurations. This enhanced convergence speed arises from the absence of an upper bound in the ReLU, facilitating larger weight updates.
Figure 2: Performance of EDLA with the sigmoid AF across various learning rates on the parity check task. The left column depicts the accuracy of the EDLA network with the sigmoid AF as a function of neurons in a single hidden layer for different learning rates (LR = 0.1, 1.0, 10.0). The accuracies are averaged from 19900 to 20000 epochs. The right column shows the number of epochs required to reach an accuracy of 0.9 as a function of neurons in a hidden layer for the same learning rates. Each row of plots corresponds to a different bit length (n_bit, ranging from 2 to 5), for the parity check task. The results are averaged over ten independent trials.
Figure 3: Accuracy and convergence characteristics of the EDLA network with the sigmoid AF across various network depths (1, 2, 4, and 8 hidden layers) at a fixed learning rate of 1.0. The left column shows the average classification accuracy as a function of the neurons per hidden layer. The accuracies are average from 19900 to 20000 epochs. The right column presents the number of epochs required to reach an accuracy of 0.9 as a function of neurons per hidden layer. Each row of plots corresponds to different bit lengths (n_bit, ranging from 2 to 5). The results are averaged over ten independent trials.
Figure 4: Performance of EDLA with the ReLU AF across various learning rates in parity check task. The left column shows the accuracy of the EDLA network as a function of neurons in a single hidden layer for different learning rates (LR = 0.01, 0.1, 1.0). The accuracies are average from 19900 to 20000 epochs. The right column shows the number of epochs required to achieve an accuracy of 0.9 against the number of neurons in the hidden layer for the same learning rates. Each row of plots corresponds to a different bit number (n_bit), ranging from 2 to 5, for the parity check task. Missing data points in the left column indicate that the accuracy is excessively low. Missing data points in the right column indicate that the accuracy does not reach 0.9. The results are averaged over ten independent trials.
Figure 5: Accuracy and convergence characteristics of the EDLA network with ReLU AF across various network depths (1, 2, 4, and 8 hidden layers), using a fixed learning rate of 1.0. The left column shows average classification accuracy as a function of neurons per hidden layer. The accuracies are average from 19900 to 20000 epochs. The right column presents the number of epochs required to reach an accuracy of 0.9 as a function of the number of neurons per hidden layer. Each row of plots corresponds to different bit lengths (n_bit), ranging from 2 to 5. The results are averaged over ten independent trials.
4.2 Regression Tasks
The EDLA network is evaluated on regression tasks using three benchmark datasets: Airfoil Self-Noise, Concrete Compressive Strength, and Energy Efficiency. The Airfoil Self-Noise dataset comprises 1503 samples with five features and a target variable representing the sound pressure level. The Concrete Compressive Strength dataset includes 1030 samples with eight features and a target variable denoting the compressive strength of concrete. The Energy Efficiency dataset comprises 768 samples with eight features and two target variables: heating and cooling loads. In this experiment, only the heating load is employed as the target variable, and the cooling load is not considered. The training and test datasets are randomly split into 80% and 20%, respectively, for all datasets.
The EDLA network consists of an input, hidden, and output layer. The neurons in the hidden layers use the sigmoid or ReLU AFs. The neurons in the output layer use the linear AF (i.e., no AF). The training parameters are carefully selected for better performance, with a learning rate of 0.001 for the sigmoid AF and 0.0001 for the ReLU AF. Training is conducted for 500 epochs employing a mini-batch size of 64 samples.
Figure 6 illustrates the Mean Absolute Error (MAE) performance of the EDLA network across varying numbers of neurons (8, 16, 32, 64, 128, 256, 512, and 1024) and different layer configurations (1, 2, and 4 layers). The results indicate that EDLA networks using sigmoid AF with one and two layers achieve efficient performance across almost all neuron-count configurations for all datasets. The EDLA network, which has a single hidden layer using the ReLU AF, also achieves efficient performance. This improves performance as the number of neurons in the hidden layers increases. Additionally, increasing the network depth does not yield performance enhancements and exhibits deterioration compared to shallower network architectures. This deterioration will be caused by the instability of the EDLA network due to the increase in both neuron count and network depth. These observations collectively suggest that, while the EDLA approach is adequately capable of regression tasks, augmenting the number of neurons and network depth does not necessarily improve performance.
Figure 6: Performance of the EDLA network on regression datasets across varying layer configurations. The left column shows the Mean Absolute Error (MAE) for sigmoid AF, while the right column displays the MAE for the ReLU AF. Both sets of results are plotted as a function of the number of neurons per hidden layer for 1, 2, and 4 hidden layers. Each row corresponds to the Airfoil Self-Noise dataset (top row), Concrete Compressive Strength dataset (middle row), and Energy Efficiency dataset (bottom row). Missing data points in the figure signify that the network output becomes excessively large or yields a not-a-number (NaN) value. The results are averaged over four independent trials.
4.3 Image Classification Tasks
To evaluate the performance of the EDLA network in image classification tasks, the experiments are conducted using the handwritten digits dataset (Digits), MNIST, and Fashion MNIST datasets. The Digits used consists of 5620 grayscale images with a resolution of $8\times 8$ pixels. Because the Digits dataset is not divided into training and test datasets, 20 % of the entire dataset is randomly selected and used as test data. The MNIST comprises 60,000 training and 10,000 testing grayscale images of handwritten digits, each with a resolution of $28\times 28$ pixels. Similarly, the Fashion-MNIST includes 60,000 training and 10,000 testing grayscale images of fashion articles at a resolution of $28\times 28$ pixels.
The EDLA network consists of an input layer, a hidden layer, and an output layer. The neurons in the input and hidden layers utilize the sigmoid or ReLU AFs. The output layer employs the sigmoid AF. To assess its performance across various configurations, the number of neurons in each hidden layer is varied systematically. The learning rates when using the sigmoid and ReLU AFs are 1.0 and 0.1, respectively. The network is trained for 500 epochs. We employed the minibatch method and set the minibatch size to 128 samples.
Figure 7 illustrates the performance of the EDLA network using the sigmoid AF for various numbers of neurons in each hidden layer. Across all datasets, the classification performance generally improved as the number of neurons in the hidden layers increased. Specifically, for the Digits and MNIST datasets, near-perfect accuracy is rapidly achieved, even with a relatively small number of neurons. For the Fashion MNIST datasets, while high accuracy is attained, a larger number of neurons contributed to an incrementally improved performance. However, the depth of the EDLA network does not improve the performance, but worsens the performance. In terms of the reach epoch, for all configurations, the EDLA network shows a sufficiently fast reach of 0.8 accuracy. While deeper networks will lead to a slower reach, the number of neurons in each hidden layer will lead to a faster reach.
Similarly, Figure 8 shows the performance of the EDLA network using the ReLU AF across varying numbers of neurons in each hidden layer. The results reveal that the EDLA network with the ReLU AF achieves a performance comparable to that of a sigmoid-based network. However, for the ReLU AF, deeper networks lead to worse performance. Furthermore, for a deeper network, the number of neurons in each hidden layer leads to worse performance. The reach epoch for the ReLU AF decreases with the number of neurons in each hidden layer. However, the reach epoch for deeper networks will lead to slower reach.
Figure 7: Performance of EDLA with the sigmoid AF on image datasets across layer configurations. The plots in the left column show the accuracy of the EDLA network with the sigmoid AF as a function of the number of neurons in each hidden layer for different layer configurations (1, 2, and 4 layers) across three datasets: digits, MNIST, and fashion_MNIST. The plots in the right column illustrate the number of epochs required to achieve an accuracy of 0.8 against the number of neurons in each hidden layer for the same layer configurations. The results are averaged over four independent trials.
Figure 8: Performance of EDLA with the ReLU AF on image datasets across layer configurations. The plots in the left column show the accuracy of the EDLA network with the sigmoid AF as a function of the number of neurons in each hidden layer for different layer configurations (1, 2, and 4 layers) across three datasets: digits, MNIST, and fashion_MNIST. The plots in the right column illustrate the number of epochs required to achieve an accuracy of 0.8 against the number of neurons in each hidden layer for the same layer configurations. The results are averaged over four independent trials.
4.4 Analysis of Internal Representations
In this subsection, we investigate the internal representations learned by the EDLA network by analyzing the synaptic weights in the hidden layers. The experiments use an EDLA network with a single hidden layer containing two neurons with sigmoid AFs. The network is trained on the Digits dataset over 500 epochs.
Figure 9 shows a heatmap of the synaptic weights connecting the input layer to the hidden layer within the EDLA network, specifically trained to classify the digit “zero” from handwritten digits. The heatmap reveals that the network learned meaningful features from the input data. Notably, the neurons in the hidden layer exhibit a strong response to specific patterns within the images. For instance, the weights connecting the positive neurons in the input layer to the positive neurons in the hidden layer (denoted as "P-P" connections) show high activation, shaping the digit “zero.” This result suggests that the EDLA method effectively enables neurons to capture and represent the relevant digit-specific features. Additionally, similar "zero-shaped" patterns also emerge in other synaptic connections.
Synaptic weight distributions in neural networks provide valuable insights into the learning behavior of the network and features extracted from the input data. To explore this further, we analyze the distribution of weights across the EDLA network. Figure 10 shows a histogram representing the synaptic weights within the trained EDLA network. The distribution appears to be predominantly centered around small magnitudes, indicating that most weights are relatively small. However, a distinct spread is observed with a smaller subset of synapses exhibiting large positive or negative values. This broad distribution suggests that the EDLA update unilaterally increase the absolute values of the weights while balancing the excitatory and inhibitory weights.
Figure 9: A heatmap of the synaptic weights connecting the input layer to the first hidden layer in the EDLA network is trained to classify the digit “zero” in the Digits dataset. Each subplot shows the weight patterns for individual connection patterns, labeled according to connection types: P-P (positive input to positive hidden sublayers), N-P (negative input to positive hidden sublayers), P-N (positive input to negative hidden sublayers), and N-N (negative input to negative hidden sublayers). Color gradients denote weight values, with yellow indicating strongly positive weights and purple indicating strongly negative weights.
Figure 10: Histogram of synaptic weight distribution in the EDLA Network. The horizontal axis denotes the synaptic weight values, and the vertical axis indicates the corresponding frequencies.
5 Conclusion
The Error Diffusion Learning Algorithm (EDLA) provides a biologically inspired alternative to conventional backpropagation methods for training neural networks and is characterized by the diffusion of a global error signal across network layers. This mechanism is analogous to reward-based learning processes observed in biological neural systems. Moreover, the EDLA mirrors neuromodulatory systems, such as the dopaminergic system, which projects broadly across cortical and subcortical areas, disseminating their signals globally (Roelfsema and Holtmaat, 2018). In contrast to standard MLPs, the EDLA incorporates structured interactions between neurons with excitatory and inhibitory synapses. This reflects the biologically observed balance between these types of synapses.
However, EDLA is inherently restricted by its single-output architecture, which limits its applicability in multiclass classification tasks. A practical solution involves parallelizing multiple EDLA networks, although this approach increases computational costs. This parallelized architecture mirrors the modular structure of biological neural systems in which specialized circuits process distinct features or tasks independently. In such a configuration, each EDLA network separately optimizes its own output according to a unique global error signal to that network. This modular structure aligns with distributed processing mechanisms observed in biological neural networks.
The EDLA demonstrates efficacy across various tasks, including parity checks, regression, and image classification. Experimental evaluations indicate that the EDLA achieves a satisfactory performance across various datasets. Notably, the ReLU AF achieves a performance comparable to that of the sigmoid function originally proposed by Kaneko (Kaneko, 1999), where the ReLU often converges more quickly. However, because the ReLU lacks an upper bound, it can produce huge and potentially unstable weight updates under the Hebbian-like learning dynamics employed by the EDLA. Moreover, while increasing the number of neurons in a hidden layer improves the performance and accelerates convergence, excessive network depth or an overly large number of hidden neurons may degrade the performance because of unstable weight updates. A reduction in the learning rate can mitigate these instabilities.
In contrast to backpropagation, which requires error signals to be propagated backward through each intermediate layer, EDLA relies solely on a single global error signal combined with local neural activities. This approach presents potential advantages for developing more biologically plausible neural network training techniques. Thus, the EDLA will represent a significant step toward bridging the gap between biological realism and artificial neural network methodologies.
Acknowledgments
The author gratefully acknowledges the assistance of large language models in improving the grammar and style of this work.
Appendix
Appendix A Multilayer Perceptron and Backpropagation
A multilayer perceptron (MLP) is a class of feedforward artificial neural networks characterized by multiple interconnected neuron layers: an input layer, one or more hidden layers, and an output layer. Typically, each neuron within one layer is fully connected to all neurons in the subsequent layer.
The input to an MLP is represented by the vector $\mathbf{x}=[x_{0},x_{1},\ldots,x_{N}]^{\mathrm{T}}\in\mathbb{R}^{N+1}$, where $x_{0}$ commonly serves as a bias term (usually set as $x_{0}=1$) and $N$ denotes the number of input features. This input vector is received by the input layer and subsequently propagated forward to the first hidden layer.
For neuron $j$ in the first hidden layer $(l=1)$, the activation $a_{j}^{(1)}$ is calculated as the weighted sum of its inputs:
$a_{j}^{(1)}=\mathbf{w}_{j}^{(1)\mathrm{T}}\mathbf{x}$ (17)
where $\mathbf{w}_{j}^{(1)}=[w_{j0}^{(1)},w_{j1}^{(1)},\ldots,w_{jN}^{(1)}]^{\mathrm{T}}$ represents the weight vector connecting the input layer to neuron $j$ in the first hidden layer $(l)$. Here, $w_{ji}^{(1)}$ denotes the weight connecting from the neuron $i$ in the input layer to the neuron $j$ in the first hidden layer. The output of neuron $j$ is then obtained by applying the activation function $g(\cdot)$:
$z_{j}^{(1)}=g(a_{j}^{(1)}).$ (18)
For subsequent hidden layers $(l=2,...,L)$, the activation $a_{j}^{(l)}$ for neuron $j$ in layer $l$ is computed as follows:
$a_{j}^{(l)}=\mathbf{w}_{j}^{(l)\mathrm{T}}\mathbf{z}^{(l-1)},$ (19)
where $\mathbf{z}^{(l)}=[z_{0}^{(l)},z_{1}^{(l)},\ldots,z_{M^{(l)}}^{(l)}]^{\mathrm{T}}\in\mathbb{R}^{M^{(l)}+1}$ is the output vector of layer $l$, $M^{(l)}$ is the number of neurons in this layer, and $z_{0}^{(l)}$ corresponds to the output of a bias neuron fixed at one. $w_{ji}^{(l)}$ denotes the weight connecting neuron $i$ in layer $l-1$ to neuron $j$ in layer $(l)$. The output for neuron $j$ is then given by
$z_{j}^{(l)}=g(a_{j}^{(l)}).$ (20)
Finally, the output for neuron $k$ in the output layer $(L+1)$ is similarly computed as:
$y_{k}=f(a_{k}^{(L+1)})=f(\mathbf{w}_{k}^{(L+1)\mathrm{T}}\mathbf{z}^{(L)})$ (21)
where $f(\cdot)$ is the activation function for the output neurons, and $\mathbf{w}_{k}^{(L+1)}$ is the weight connecting the neurons in the final hidden layer $(L)$ to output neuron $k$.
Training an MLP seeks the optimal weights $\mathbf{w}$ that minimize a predetermined loss function $E(\mathbf{w})$, such as mean squared error or cross-entropy loss. Gradient descent is a widely used optimization method, in which weights are iteratively updated in the direction that locally reduces the value of $E(\mathbf{w})$. To efficiently compute gradients necessary for this optimization, the backpropagation algorithm employs the chain rule, propagating the error signals sequentially from the output layer backward through each hidden layer to the input layer.
Consider an output neuron $k$ in the output layer $(L+1)$. Using the chain rule, the partial derivative of the loss function $E(\mathbf{w})$ with respect to the weight $w_{kj}^{(L+1)}$ connecting neuron $j$ in the last hidden layer $(L)$ to neuron $k$ in the output layer is computed as follows:
$\frac{\partial E(\mathbf{w})}{\partial w_{kj}^{(L+1)}}=\frac{\partial E(\mathbf{w})}{\partial a_{k}^{(L+1)}}\frac{\partial a_{k}^{(L+1)}}{\partial w_{kj}^{(L+1)}}=\delta_{k}^{(L+1)}z_{j}^{(L)}$ (22)
where
$\delta_{k}^{(L+1)}\equiv\frac{\partial E(\mathbf{w})}{\partial a_{k}^{(L+1)}}=f^{\prime}(a_{k}^{(L+1)})\frac{\partial E(\mathbf{w})}{\partial y_{k}}.$ (23)
This quantity, often referred to as the error for output neuron $k$, is the partial derivative of the loss $E(\mathbf{w})$ with respect to the neuron’s activation $a_{k}^{(L+1)}$. By chain rule, this derivative can be factored into two components: the partial derivative of the loss function with respect to the neuron’s output, $\frac{\partial E(\mathbf{w})}{\partial y_{k}}$, and and the derivative of the activation function evaluated at $a_{k}^{(L+1)}$, namely $f^{\prime}(a_{k}^{(L+1)})$.
To calculate the error $\delta_{j}^{(l)}$ for hidden layer $(l)$, we make use of the chain rule again. The partial derivative of the error function with respect to the activation $a_{j}^{(l)}$ is:
$\frac{\partial E(\mathbf{w})}{\partial w_{ji}^{(l)}}=\frac{\partial E(\mathbf{w})}{\partial a_{j}^{(l)}}\frac{\partial a_{j}^{(l)}}{\partial w_{ji}^{(l)}}=\delta_{j}^{(l)}z_{i}^{(l-1)}.$ (24)
The error $\delta_{j}^{(l)}$ is defined as:
$\delta_{j}^{(l)}\equiv\frac{\partial E(\mathbf{w})}{\partial a_{j}^{(l)}}=g^{\prime}(a_{j}^{(l)})\sum_{k}\delta_{k}^{(l+1)}w_{kj}^{(l+1)},$ (25)
where $g^{\prime}(\cdot)$ represents the derivative of the activation function $g(\cdot)$ evaluated at $a_{j}^{(l)}$. Intuitively, the error at hidden neuron $j$ is computed based on the weighted sum of error signals from the neurons in the subsequent layer $(l+1)$, thus sequentially propagating the error signal backward through the network.
Once the partial derivatives $\frac{\partial E(\mathbf{w})}{\partial w_{ji}^{(l)}}$ are computed for all weights, the update rule in standard gradient descent is:
$w_{ji}^{(l)}\leftarrow w_{ji}^{(l)}-\eta\frac{\partial E(\mathbf{w})}{\partial w_{ji}^{(l)}},$ (26)
where $\eta$ is the learning rate. This iterative cycle—consisting of a forward pass (computation of neuron activations), a backward pass (error computation), and weight updates—continues until a predetermined convergence criterion is satisfied or a specified number of training epochs has been reached.
The gradient of the error function with respect to the weights is computed using training data. In practice, to compute the gradient, most practitioners use a method called stochastic gradient descent (SGD) (LeCun et al., 2015). In SGD, the calculation of the gradient is performed on a subset of the training data, known as a mini-batch. This approach enables efficient computation of the gradient, making it feasible to train large neural networks using extensive datasets.
Traditional MLPs have an advantage when handling network architectures requiring multiple outputs. MLPs can simultaneously optimize the weights associated with each output. This is attributed to the backpropagation mechanism in MLPs, where the weights connecting the output layer to the preceding hidden layer are specifically adjusted to minimize the error associated with each output. Concurrently, hidden-layer neurons are trained to extract and generalize pertinent features from the input data. This hierarchical optimization enables MLPs to efficiently model complex input-output relationships and successfully generalize the learned features.