### v3

# A Basic Evaluation of Neural Networks Trained with the Error Diffusion Learning Algorithm

Abstract
Background / introduction This paper presents a comprehensive formulation of Kaneko’s Error Diffusion Learning Algorithm (EDLA) and evaluates its effectiveness across parity check, regression, and image classification tasks. EDLA is a biologically inspired learning algorithm that provides an alternative to conventional backpropagation for training artificial neural networks.
Methods EDLA employs a single global error signal that diffuses across networks composed of paired positive and negative sublayers, eliminating traditional layer-wise error backpropagation. This study evaluates EDLA’s effectiveness using benchmark tasks, such as parity check, regression, and image classification, by systematically varying the neuron count, network depth, and learning rates to assess its performance comprehensively.
Results The experimental results demonstrate that EDLA achieves consistently high accuracy across multiple benchmarks, highlighting its effectiveness as a learning algorithm for neural networks. The choice of learning rate, neuron count, and network depth significantly influences EDLA’s efficiency and convergence speed. Analysis of internal network representations reveals meaningful feature extraction capabilities, and the network’s overall performance is found to be competitive with networks trained via conventional backpropagation, especially in shallow architectures.
Conclusions This study introduces EDLA, a biologically plausible alternative to traditional backpropagation previously underrecognized due to language barriers. By reformulating EDLA, systematically evaluating its performance, and presenting empirical evidence of its effectiveness, this study increases the visibility and accessibility of EDLA and contributes to biologically inspired training methodologies.
keywords:
Neural network, Learning algorithm, Error diffusion
1 Introduction
Artificial neural networks have substantially advanced numerous research domains, including image processing, signal processing, and natural language processing. The remarkable success of these networks largely stems from the backpropagation method proposed by Rumelhart et al. [1], which has become a standard method for training neural networks across diverse architectures.
The theoretical foundation of neural networks traces back to the linear threshold unit (neuron) introduced by McCulloch and Pitts [2, 3]. The fundamental mechanism of this neuron is based on the all-or-none law [4, 5] and still underlies modern neural network architectures. Rosenblatt’s development of the perceptron [6] represents another critical milestone in neural networks, marking the first realization of a trainable artificial neuron. However, because this trainable perceptron consists of only a single computational layer, it cannot inherently form nonlinear decision boundaries. Although multilayer perceptrons (MLPs) resolve this limitation, training these deeper networks has been a substantial challenge. The introduction of backpropagation has provided a solution for training MLPs. Subsequently, it has become the dominant paradigm for training various deep neural network architectures.
Despite the success of backpropagation, questions remain about its biological plausibility [7, 8]. This gap between artificial and biological neural systems has driven the development of alternative learning algorithms. One example is Kaneko’s Error Diffusion Learning Algorithm (EDLA) [9]. EDLA introduces a biologically inspired architecture that incorporates both excitatory and inhibitory synapses, reflecting biological neural systems more closely. A distinctive feature of this algorithm is its error-diffusion mechanism. In this approach, the difference between the network output and target values is diffused throughout the network during training, and has been reported to enable effective training with sigmoid activations. Unlike standard backpropagation, EDLA diffuses only a single error signal throughout the network, eliminating the need for backward propagation and error calculation for each layer. This simplification addresses one of the key problems highlighted by [8]: artificial neural networks based on backpropagation require buffering past presynaptic activities and synchronizing them with the corresponding error signals propagated sequentially from higher layers.
This study provides a comprehensive overview of the EDLA framework and evaluates its performance on several benchmark tasks including parity checks, regression, and image classification. Through these evaluations, we aim to promote the EDLA framework and facilitate the development of training methods for biologically plausible neural networks.
Portions of this text were previously published as part of a preprint (https://arxiv.org/abs/2504.14814).
2 Error Diffusion Learning Algorithm (EDLA)
The Error Diffusion Learning Algorithm (EDLA) is a novel approach to neural network learning initially proposed by Kaneko [9]. Although the original model is available only in Japanese, this section provides a comprehensive reinterpretation of EDLA.
2.1 Network Architecture
Figure 1 illustrates the neural network architecture employed by the EDLA. Like traditional MLPs, EDLA adopts a multilayer structure comprising neurons and synapses. However, EDLA incorporates two types of neurons: positive neurons (p-neurons) and negative neurons (n-neurons). The synaptic connections are categorized based on the types of neurons they link: excitatory synapses connect neurons of identical type, whereas inhibitory synapses connect neurons of different types. A key characteristic of the EDLA is that it has only one output neuron.
Figure 1: Schematic representation of the Error Diffusion Learning Algorithm (EDLA) network architecture. The network consists of multiple layers of positive neurons (white circles) and negative neurons (gray circles). The red and blue small circles at the connections indicate excitatory and inhibitory synapses, respectively.
2.2 Learning Rule and Mathematical Formulation
The input to the EDLA network is represented by the vector $\mathbf{x}=[x_{0},x_{1},\ldots,x_{N}]^{T}\in\mathbb{R}^{N+1}$, where $x_{0}$ corresponds to the bias set to one, and $N$ is the number of input features. In the EDLA architecture, each layer consists of positive and negative sublayers. The outputs from the positive and negative sublayers of the input layer are denoted by $\mathbf{x}^{p}$ and $\mathbf{x}^{n}$, respectively. For the input layer, $\mathbf{x}^{p}=\mathbf{x}^{n}=\mathbf{x}$.
For a neuron $j$ in the first hidden layer $(l=1)$, the activations for the positive and negative neurons, $a_{j}^{p,(1)}$ and $a_{j}^{n,(1)}$, are computed as follows:
$a_{j}^{p,(1)}={\mathbf{w}_{j}^{pp,(1)}}^{T}\mathbf{x}^{p}+{\mathbf{w}_{j}^{pn,(1)}}^{T}\mathbf{x}^{n},$ (1)
$a_{j}^{n,(1)}={\mathbf{w}_{j}^{np,(1)}}^{T}\mathbf{x}^{p}+{\mathbf{w}_{j}^{nn,(1)}}^{T}\mathbf{x}^{n}.$ (2)
In these equations, $\mathbf{w}_{j}^{pp,(1)}=[w_{j0}^{pp},\ldots,w_{jN}^{pp}]^{T}$ represents the vector of synaptic weights connecting the positive neurons in the input layer to positive neuron $j$ in the first hidden layer $(l=1)$. Similarly, $\mathbf{w}_{j}^{pn,(1)}$, $\mathbf{w}_{j}^{np,(1)}$, and $\mathbf{w}_{j}^{nn,(1)}$ denote the weight vectors for the other combinations of neuron types in the connections between the input and the first hidden layers.
The outputs of the positive and negative neurons $j$ in the first hidden layer are given by:
$z_{j}^{p,(1)}=g(a_{j}^{p,(1)})\quad\text{and}\quad z_{j}^{n,(1)}=g(a_{j}^{n,(1)}),$ (3)
where $g(\cdot)$ is the activation function. In subsequent hidden layers (l=2,,L)(l=2,...,L) where $L$ refers to the final hidden layer, the neuron outputs are given by
$z_{j}^{p,(l)}=g({\mathbf{w}_{j}^{pp,(l)}}^{T}\mathbf{z}^{p,(l-1)}+{\mathbf{w}_{j}^{pn,(l)}}^{T}\mathbf{z}^{n,(l-1)}),$ (4)
$z_{j}^{n,(l)}=g({\mathbf{w}_{j}^{np,(l)}}^{T}\mathbf{z}^{p,(l-1)}+{\mathbf{w}_{j}^{nn,(l)}}^{T}\mathbf{z}^{n,(l-1)}).$ (5)
The final output of the EDLA network $y$ is computed as
$y=g({\mathbf{w}^{pp,(L+1)}}^{T}\mathbf{z}^{p,(L)}+{\mathbf{w}^{pn,(L+1)}}^{T}\mathbf{z}^{n,(L)}).$ (6)
Synaptic weights within the EDLA network are constrained to maintain specific relationships, reflecting the excitatory and inhibitory nature of the connections:
$w^{pp,(l)}_{ji}\geq 0,\quad w^{pn,(l)}_{ji}\leq 0,\quad w^{np,(l)}_{ji}\leq 0,\quad w^{nn,(l)}_{ji}\geq 0.$ (7)
To reflect these constraints at the start of training, weights are initialized by sampling uniformly from $[0,1]$ for synapses connecting neurons of the same type and from [1,0][-1,0] for those connecting neurons of different types.
The objective of training an EDLA network is to identify the optimal parameter set $\mathbf{W}$ that minimizes a predefined error function $E(\mathbf{W})$, similar to traditional MLPs. Gradient descent is a widely used optimization method in which parameters are iteratively updated in a direction that locally reduces the value of $E(\mathbf{W})$. In practice, most practitioners use stochastic gradient descent (SGD) [10], which computes gradient estimates efficiently using subsets of training data (mini-batches). The EDLA also employs a method based on SGD.
Although the EDLA utilizes the principles of gradient descent, it incorporates an approximation in computing the errors used for weight updates. In an MLP trained by backpropagation, the error $\delta_{j}^{(l)}$ for neuron $j$ in hidden layer $l$ is computed as:
$\delta_{j}^{(l)}=g^{\prime}(a_{j}^{(l)})\sum_{k}w_{jk}^{(l+1)}\delta_{k}^{(l+1)},$ (8)
where $g^{\prime}(a_{j}^{(l)})$ is the derivative of the activation function. In contrast, the EDLA approximates these errors using a single global error signal $d$, defined as
$d=-\frac{\partial E(\mathbf{W})}{\partial y}.$ (9)
This global error is uniformly diffused throughout the network. Specifically, the approximated errors, $\tilde{\delta}^{pp}_{ji}$, $\tilde{\delta}^{np}_{ji}$, $\tilde{\delta}^{nn}_{ji}$, and $\tilde{\delta}^{pn}_{ji}$ in the EDLA, are expressed as:
$\tilde{\delta}^{pp,(l)}_{ji}=dg^{\prime}(a^{p,(l)}_{j}),\tilde{\delta}^{np,(l)}_{ji}=dg^{\prime}(a^{n,(l)}_{j})$, (10)
$\tilde{\delta}^{nn,(l)}_{ji}=dg^{\prime}(a^{n,(l)}_{j}),\tilde{\delta}^{pn,(l)}_{ji}=dg^{\prime}(a^{p,(l)}_{j})$. (11)
[…]
The weight updates in EDLA differ depending on the sign of the global error $d$. When $d>0$, the weights originating from the positive neurons $w^{pp,(l)}_{ji}$ and $w^{np,(l)}_{ji}$ are updated as follows:
$w^{pp,(l)}_{ji}\leftarrow w^{pp,(l)}_{ji}+\eta\tilde{\delta}^{pp,(l)}_{ji}z^{p,(l-1)}_{i}=w^{pp,(l)}_{ji}+\eta dg^{\prime}(a^{p,(l)}_{j})z^{p,(l-1)}_{i}$, (12)
$w^{np,(l)}_{ji}\leftarrow w^{np,(l)}_{ji}-\eta\tilde{\delta}^{np,(l)}_{ji}z^{p,(l-1)}_{i}=w^{np,(l)}_{ji}-\eta dg^{\prime}(a^{n,(l)}_{j})z^{p,(l-1)}_{i}$. (13)
Conversely, when $d<0$, the weights originating from the negative neurons, $w^{nn,(l)}_{ji}$ and $w^{pn,(l)}_{ji}$ are updated as:
$w^{nn,(l)}_{ji}\leftarrow w^{nn,(l)}_{ji}-\eta\tilde{\delta}^{nn,(l)}_{ji}z^{n,(l-1)}_{i}=w^{nn,(l)}_{ji}-\eta dg^{\prime}(a^{n,(l)}_{j})z^{n,(l-1)}_{i}$, (14)
$w^{pn,(l)}_{ji}\leftarrow w^{pn,(l)}_{ji}+\eta\tilde{\delta}^{pn,(l)}_{ji}z^{n,(l-1)}_{i}=w^{pn,(l)}_{ji}+\eta dg^{\prime}(a^{p,(l)}_{j})z^{n,(l-1)}_{i}$. (15)
where $\eta$ denotes the learning rate. Thus, the global error signal directly determines whether weights from positive or negative neurons are adjusted during training.
The behavior of the EDLA becomes intuitively clear when employing the sigmoid activation function $g(a)=1/(1+\exp(-a))$ along with the squared error loss function defined as $E(\mathbf{w})=\frac{1}{2}(y-t)^{2}$. The global error signal $d$ is represented as follows:
$d=-\frac{\partial E(\mathbf{w})}{\partial y}=-\frac{\partial}{\partial y}\left[\frac{1}{2}(y-t)^{2}\right]=t-y.$ (16)
When the network output is smaller than the target value ($d>0$), the absolute values of the weights originating from positive neurons increase due to the properties $g^{\prime}(a)>0$ and $z>0$. Such weight adjustments amplify the outputs of positive neurons in subsequent layers while simultaneously suppressing the outputs of negative neurons in those layers. Consequently, these modifications reduce the discrepancy between the output and target value of the network. Conversely, when the network output exceeds the target value ($d<0$), the absolute values of the weights originating from negative neurons increase. These weight adjustments decrease the outputs of positive neurons in subsequent layers while simultaneously increasing the outputs of negative neurons. Thus, regardless of the sign of the global error signal, EDLA systematically reduces the difference between the network output and target values through successive weight updates. Importantly, this characteristic learning behavior of EDLA is maintained as long as $g^{\prime}(a)>0$, $z>0$, and the sign of the derivative of the loss function properly encodes whether the network output is below or above the target value. The same squared error loss function was used in all experiments conducted in this study.
For the EDLA, the activation function is typically chosen as the sigmoid function. It constrains the output of the neuron between 0 and 1 and its derivative between 0 and 0.25. By limiting both the neuron outputs and their derivatives, the sigmoid activation function provides stability to the training dynamics, avoiding large weight updates and facilitating stable convergence.
2.3 Biological Interpretation
[…]
Moreover, EDLA employs a global error-diffusion mechanism in which the error signal $d$ is broadcast across the network. This diffusion is conceptually analogous to biological neuromodulation and closely parallels reward-prediction-error (RPE) signals observed in biological systems, which guide synaptic plasticity via diffuse neuromodulatory pathways (e.g., dopamine) [12], thereby mirroring how neuromodulators broadly regulate synaptic change.
Unlike traditional Hebbian learning—which is typically unsupervised and driven solely by local activity correlations—EDLA integrates a supervised component through the global error signal $d$, modulating synaptic changes according to network performance relative to a target. Conceptually, EDLA aligns with neo-Hebbian three-factor rules in which synaptic change depends on (i) a presynaptic factor, (ii) a postsynaptic factor, and (iii) a systems-level modulatory factor [13, 14]. In EDLA the error $d$ plays this third role: weight updates depend on presynaptic activity $z^{p,(l-1)}_{i}$, a postsynaptic state $g^{\prime}(a^{p,(l)}_{j})$, and the global error signal $d$.
In canonical neuroscientific three-factor formulations, pre- and postsynaptic coincidence first sets a synapse-specific eligibility trace (or tag), and a later modulatory signal (e.g., dopamine/RPE) converts that trace into an actual weight update. Eligibility traces bridge the temporal gap between neural activity and delayed reward signals, which is particularly important in spiking networks where modulatory signals often arrive with delays relative to the activity that made a synapse “eligible” [14]. By contrast, EDLA operates in an artificial-network regime where the global error signal is computed and broadcast immediately at each training step. Accordingly, EDLA may be viewed as a limiting case of three-factor rules in which eligibility traces have effectively negligible lifetime and modulatory delays are absent; this approximation is reasonable for supervised tasks with immediate error availability but limits applicability in scenarios with delayed modulatory feedback.
2.4 Limitation
The inherent simplicity and approximation underlying EDLA introduce significant constraints, particularly concerning the network’s capability to handle multiple outputs effectively. In the EDLA framework, the global error signal $d$ is uniformly diffused throughout the network architecture. This global error diffusion inherently directs the optimization process toward minimizing the error associated with a single-output node. Consequently, when dealing with single-output scenarios, the EDLA can efficiently and accurately specialize the network weights for a specific output. However, when an EDLA network contains multiple outputs, each output generates its own global error signal, which propagates throughout the network. Such simultaneous diffusion of distinct error signals leads to conflicting weight adjustments, as the network simultaneously attempts to minimize errors for multiple outputs. These conflicting updates inhibit the network’s ability to extract and generalize meaningful features from the input data, thereby undermining learning efficacy.
2.5 Extending EDLA to Multiple Outputs
The inherent single-output constraint of the EDLA represents a significant limitation when addressing tasks that require multiple outputs. A simple approach to adapting EDLA to overcome this limitation involves constructing multiple EDLA networks arranged in parallel. This parallelized architecture addresses a task that requires $K$ outputs by employing $K$ independent EDLA networks. Each individual network retains the fundamental EDLA architecture and operates autonomously. Input vector $\mathbf{x}$ is simultaneously provided to all networks, ensuring that they process the same input data. Each network independently computes output $y_{k}$, where $k\in\{1,\ldots,K\}$. In classification tasks, if the output $y_{k}$ is the maximum value among the outputs of the EDLA networks, $k$ is regarded as the overall output label.
The weights of each individual EDLA network are independently optimized using the standard EDLA update rule. Consequently, this parallel architecture retains the capability, simplicity, and core principles inherent to single-output EDLA models, effectively extending EDLA’s capability to handle multi-output tasks. However, this parallelization approach increases computational costs because the total number of EDLA networks scales linearly with the required number of outputs. This trade-off highlights the challenge of maintaining the computational simplicity of single-output EDLA networks while contending with the additional resource demands necessary for multi-output learning.
Employing multiple single-output networks in an EDLA framework may seem computationally inefficient compared with traditional multi-output MLP architectures. Nevertheless, this parallelization method is reminiscent of modular structures, which are organizational principles observed in biological neural systems, particularly in the brain [15]. Mountcastle’s early neurophysiological research [16] has demonstrated that neurons within the primate somatosensory cortex are organized into vertically aligned columns, each dedicated to a particular modality or receptive field. Additionally, neurons in the primary visual cortex are organized into modular structures known as hypercolumns, each specializing in processing specific features of visual stimuli such as orientation [17]. Each hypercolumn functions semi-independently, collectively contributing to overall visual perception by handling distinct components of the visual scene. Similarly, the inferior temporal cortex exhibits clusters of neurons that are selectively responsive to complex visual features including distinct object shapes [18, 19]. This clustered arrangement enables the efficient processing of diverse visual elements while simultaneously facilitating effective integration across features and categories. Another notable example is the barrel cortex of rodents, in which neural clusters (called barrels) are dedicated to independently processing tactile inputs from individual whiskers [20, 21]. Each barrel independently processes sensory information corresponding to its associated whisker, and the combined neural activity across the barrels generates a comprehensive tactile representation. These modular and distributed neural processing strategies enable the brain to handle complex and heterogeneous tasks efficiently while preserving both flexibility and robustness. Although the parallelized architecture of EDLA, which consists of multiple single-output networks dedicated to specific outputs, may be computationally less efficient than traditional multi-output MLPs, it reflects these biological modular architectures.
The increased computational cost of the parallelized EDLA is an important practical consideration. However, the increase may be less problematic for brain systems or neuromorphic computing applications. In such systems, neurons and synapses are implemented as individually operating, massively parallel [22, 23]. As a result, increases in neuron and synapse counts do not necessarily translate into proportional increases in wall-clock computation time. Some areas of the brain discussed above, such as the visual cortex and the barrel cortex, employ modular architectures that process complex sensory inputs in real-time. Indeed, researchers have achieved real-time simulation using massively parallel systems [22]. Therefore, while the parallelized EDLA entails a trade-off between biological plausibility and computational efficiency on standard computer systems, it may be particularly well suited to neuromorphic implementations, brain-inspired computing systems, or other massively parallel platforms. However, the increase in the number of unit circuits for neuromorphic systems and the memory cost for parallel computing remain a concern, as each EDLA network requires its own set of weights and parameters.
2.6 Number of parameters in EDLA
For an EDLA network with $L$ hidden layers, $n_{\mathrm{hid}}$ positive neurons and $n_{\mathrm{hid}}$ negative neurons per layer, input size $n_{\mathrm{in}}$, and $n_{\mathrm{out}}$ outputs (implemented as $n_{\mathrm{out}}$ independent single-output EDLA networks), the total number of parameters, $N_{\mathrm{EDLA}}$, is
$N_{\mathrm{EDLA}}$ $=\big(2n_{\mathrm{in}}(2n_{\mathrm{hid}})+(L-1)(2n_{\mathrm{hid}})^{2}+2n_{\mathrm{hid}}\big)\,n_{\mathrm{out}}$ (17)
$=2n_{\mathrm{out}}\,n_{\mathrm{hid}}\big(2n_{\mathrm{in}}+2(L-1)n_{\mathrm{hid}}+1\big).$ (18)
Here the first term, $2n_{\mathrm{in}}(2n_{\mathrm{hid}})$, corresponds to the weights connecting the input layer to the first hidden layer; the second term, $(L-1)(2n_{\mathrm{hid}})^{2}$, accounts for the weights between hidden layers; and the third term, $2n_{\mathrm{hid}}$, corresponds to the weights from the last hidden layer to a single output (there are $n_{\mathrm{hid}}$ positive and $n_{\mathrm{hid}}$ negative neurons, hence $2n_{\mathrm{hid}}$). Multiplying by $n_{\mathrm{out}}$ yields the total number of parameters across all independent single-output EDLA networks. Bias terms are omitted in this theoretical analysis for simplicity. The exact parameter counts reported in Table 1, which are used for experimental comparison, include bias terms.
In contrast, the parameter count for a standard fully connected MLP with $L$ hidden layers, $m_{\mathrm{hid}}$ neurons per hidden layer, input size $n_{\mathrm{in}}$, and $n_{\mathrm{out}}$ outputs, $N_{\mathrm{MLP}}$, is
$N_{\mathrm{MLP}}$ $=n_{\mathrm{in}}m_{\mathrm{hid}}+(L-1)m_{\mathrm{hid}}^{2}+m_{\mathrm{hid}}n_{\mathrm{out}}$ (19)
$=m_{\mathrm{hid}}\big(n_{\mathrm{in}}+(L-1)m_{\mathrm{hid}}+n_{\mathrm{out}}\big),$ (20)
again omitting bias terms. Unlike EDLA, the parameter count of a traditional MLP does not scale strongly with $n_{\mathrm{out}}$ (it appears only in the final layer term).
To compare parameter efficiency, consider the case $n_{\mathrm{hid}}=m_{\mathrm{hid}}$ and assume the network is sufficiently deep and wide so that hidden-to-hidden weights dominate. When $2(L-1)n_{\mathrm{hid}}\gg 2n_{\mathrm{in}}+1$, the EDLA parameter count can be approximated as
$N_{\mathrm{EDLA}}\approx 4\,n_{\mathrm{out}}\,n_{\mathrm{hid}}^{2}(L-1).$ (21)
If $n_{\mathrm{hid}}=m_{\mathrm{hid}}$ and $(L-1)m_{\mathrm{hid}}\gg n_{\mathrm{in}}+n_{\mathrm{out}}$, then
$N_{\mathrm{MLP}}\approx n_{\mathrm{hid}}^{2}(L-1).$ (22)
Thus, in this limit,
$\frac{N_{\mathrm{EDLA}}}{N_{\mathrm{MLP}}}\approx 4\,n_{\mathrm{out}}.$ (23)
An EDLA implemented as $n_{\mathrm{out}}$ independent single-output networks can therefore require roughly $4\,n_{\mathrm{out}}$ times more parameters than a comparable MLP with the same hidden-layer size and depth. This substantial increase in the parameter count can raise memory usage and computational cost during training and inference, particularly for tasks with many outputs. The computational costs may, however, be partially mitigated by parallel implementations or neuromorphic hardware, as noted in the previous subsection. If a multi-output EDLA architecture were developed that shares computations across outputs in future work, it could substantially reduce the parameter overhead compared with the current parallelized approach.
2.7 Stabilization by RMS Normalization
When the activation function has no upper bound (for example, ReLU), individual pre-activation values and the corresponding outputs can grow without bound. Consequently, the magnitude of weight updates depends directly on the activation of the preceding layer. In particular, large pre-activations under ReLU can produce very large updates and thereby destabilize the learning dynamics. To mitigate this problem, Root Mean Square (RMS) normalization [24] is applied to the pre-activations. Normalization techniques such as RMS, Layer Normalization [25], and batch normalization [26] are commonly applied to pre-activations in artificial neural networks.
The RMS is defined as
$\mathrm{RMS}=\sqrt{\frac{1}{2n_{\mathrm{hid}}}\sum_{j=1}^{n_{\mathrm{hid}}}\left((a^{p,(l)}_{j})^{2}+(a^{n,(l)}_{j})^{2}\right)}$ (24)
where $n_{\mathrm{hid}}$ is the number of neurons in each sublayer (so that $2n_{\mathrm{hid}}$ is the total number of hidden neurons in the full layer). The normalized pre-activation $\tilde{a}^{p,(l)}_{j}$ is then given by
$\tilde{a}^{p,(l)}_{j}=\frac{a^{p,(l)}_{j}}{\mathrm{RMS}}\gamma.$ (25)
An analogous expression holds for $\tilde{a}^{n,(l)}_{j}$. The positive scalar $\gamma$ controls the overall magnitude of the normalized pre-activations. In this study, $\gamma$ is fixed to 1.0 because learning $\gamma$ is outside the scope of this work and because $\gamma$ is here treated as a biologically motivated gain parameter. A fixed $\gamma$ scales the pre-activation vector of each layer without changing its direction. RMS normalization is applied only between hidden layers; it is not applied between the input layer and the first hidden layer, nor between the final hidden layer and the output layer.
Incorporating RMS normalization, the weight update rule for a connection from a positive presynaptic neuron to a positive postsynaptic neuron is derived as follows:
$dw^{pp,(l)}_{ji}$ $=\eta dg^{\prime}(\tilde{a}^{p,(l)}_{j})\frac{\partial\tilde{a}^{p,(l)}_{j}}{\partial a^{p,(l)}_{j}}z^{p,(l-1)}_{i}$ (26)
$=\eta dg^{\prime}(\tilde{a}^{p,(l)}_{j})\frac{1}{\mathrm{RMS}}\left(1-\frac{{a^{p,(l)}_{j}}^{2}}{2n_{\mathrm{hid}}\mathrm{RMS}^{2}}\right)z^{p,(l-1)}_{i}$ (27)
The term $1-\frac{{a^{p,(l)}_{j}}^{2}}{2n_{\mathrm{hid}}\mathrm{RMS}^{2}}$ is positive because $2n_{\mathrm{hid}}\mathrm{RMS}^{2}$ is the sum of all squared pre-activations in the layer, which is necessarily greater than any single squared pre-activation $(a^{p,(l)}_{j})^{2}$. Therefore, the sign of the weight update $dw^{pp,(l)}_{ji}$ remains determined by the sign of the global error signal $d$, preserving the fundamental learning mechanism of the EDLA. The update rules for other weight types are derived in an analogous manner.
RMS normalization can be interpreted in the context of biological information processing, relating to mechanisms like gain control and divisive normalization. Gain control [27, 28] modulates neuronal sensitivity according to the overall activity level of a population, thereby helping to maintain firing rates within a stable dynamic range and avoiding saturation. Divisive normalization [29, 30] divides a neuron’s response by a pooled measure of population activity, preventing any single unit from dominating the representation. RMS normalization performs a related operation by scaling pre-activations in proportion to the population RMS; this reduces the magnitude of excessively large updates and thus stabilizes learning.
RMS normalization is selected over Layer Normalization due to a key property of the EDLA architecture. In EDLA, non-negative activation functions such as the sigmoid or ReLU are frequently employed. Layer Normalization standardizes activations to have zero mean and unit variance, a process that involves re-centering the data. This re-centering inevitably shifts many pre-activation values into the negative range. In particular, when using the ReLU activation function, this would cause a significant portion of neurons to output zero, effectively silencing them and impeding the weight updates. RMS normalization avoids this issue by rescaling activations without the re-centering step.
Note that RMS normalization is applied only in the experiments reported in Section 4.5; it is omitted in the other experiments described in this manuscript.
3 Experimental Setup
The neural networks are implemented in Python using NumPy for linear algebra operations and PyTorch for neural networks. The source code is available on GitHub at https://github.com/KazuhisaFujita/EDLA for reproducibility and further extension of this work.
4 Experimental Results
4.1 Parity Check Task
To evaluate the efficacy of the Error Diffusion Learning Algorithm (EDLA), we initially assessed its performance on a parity check task, a classic benchmark in neural network learning. This task involves determining whether the number of ones in a binary input vector is even or odd. The parity check task is a two-class classification problem. The EDLA network consists of input, hidden, and output layers. The neurons in the hidden layers use the sigmoid or ReLU activation functions (AFs). The network has a single-output neuron that employs the sigmoid AF, producing a value between 0 and 1. The output is interpreted as the probability of the input vector containing an odd number of ones. Thus, the output neuron is expected to yield a value above 0.5 for odd counts and below 0.5 for even counts. The total number of possible input combinations is $2^{n_{\mathrm{bit}}}$, where $n_{\mathrm{bit}}$ denotes the number of bits in the input vector. In this experiment, $n_{\mathrm{bit}}=2,3,4,5$ with each dataset including all possible combinations for a given $n_{\text{bit}}$. A mini-batch size of 4 is consistently employed across all configurations.
Figure 2 illustrates the performance of the EDLA network with a single hidden layer and the sigmoid AF across various learning rates (0.1, 1.0, 10.0) and neuron counts per layer (8, 16, 32, 64, 128). The results demonstrate that the EDLA network achieves high accuracy on the parity check task. The network attains perfect accuracy across all cases except for configurations with eight neurons when $n_{\mathrm{bit}}=5$. The epochs required to reach an accuracy of 0.9 decrease as the number of neurons increase. For learning rates of 10 and 1.0, the convergence time (reach epoch) is shorter than that for a learning rate of 0.1. Furthermore, for $n_{\mathrm{bit}}\in\{3,5\}$, the reach epoch at a learning rate of 1.0 is nearly identical to that at a learning rate of 10. For $n_{\mathrm{bit}}=2$, the reach epoch for a learning rate of 1.0 is larger than that at a learning rate of 10, but for $n_{\mathrm{bit}}=4$ the reach epoch at a learning rate of 1.0 is smaller than that at a learning rate of 10. These results suggest that learning rates of 10 and 1.0 enable efficient training.
Figure 3 presents an analysis of the performance of the EDLA network using the sigmoid AF across various hidden layers (1, 2, 4, and 8 hidden layers) and neuron counts per layer with a fixed learning rate of 1.0. The results reveal that the accuracy consistently approaches or reaches 1.0 across different bit numbers ($n_{\mathrm{bit}}$), regardless of the number of hidden layers. However, the number of epochs required to achieve 0.9 accuracy reveals more nuanced dependencies on the network depth and neuron count. Specifically, for $n_{\mathrm{bit}}=2$, shallow networks (one layer) exhibit a relatively low reach epoch for smaller neuron numbers, but the reach epoch increases for larger neuron counts. The deeper networks do not exhibit significant reach epoch reductions for $n_{\mathrm{bit}}=2$. Conversely, for $n_{\mathrm{bit}}\in\{3,4\}$ with deeper architectures (four and eight layers), the reach epoch increases with the number of neurons. For $n_{\mathrm{bit}}=5$, the reach epoch decreases notably with the neuron count and stabilizes for neuron counts of 64 and above. These results indicate that EDLA networks with the sigmoid AF can consistently achieve high accuracy on the parity check task, with optimal performance depending on the complexity of the task and the chosen network configuration.
Figure 4 shows the performance of the EDLA network with a single hidden layer and the ReLU AF across various learning rates (0.01, 0.1, 1.0) and neuron counts of the hidden layer. For learning rates of 0.01 and 0.1, the network achieves high accuracy on the parity check task, with the number of epochs required to reach an accuracy of 0.9, decreasing as the number of neurons increases. For a learning rate of 0.01, the network exhibits a slower convergence speed than a learning rate of 0.1, particularly for smaller neuron counts. For a learning rate of 1.0, the performance is worse than that for the other learning rates. In particular, for a learning rate of 1.0 and $n_{\mathrm{bit}}\in\{4,5\}$, the network fails to achieve an accuracy of 0.9 even with a large number of neurons. This shows that a learning rate of 1.0, which is too high for the EDLA network with the ReLU, leads to instability in training and poor performance.
Figure 5 shows the accuracy and convergence speed of the EDLA network with the ReLU across various neuron counts and hidden layer depths for a learning rate of 0.1. Both the accuracy and convergence speed improve with increasing neuron counts in the ReLU network. However, this improvement in the convergence speed does not continue indefinitely and eventually plateaus. These results indicate a performance comparable to that of sigmoid-based networks, with a consistently faster convergence observed for ReLU-based configurations. This enhanced convergence speed arises from the absence of an upper bound in the ReLU, facilitating larger weight updates.
Figure 2: Performance of EDLA with the sigmoid AF across various learning rates on the parity check task. The left column depicts the accuracy of the EDLA network with the sigmoid AF as a function of neurons in a single hidden layer for different learning rates (LR = 0.1, 1.0, 10.0). The accuracies are averaged from 19900 to 20000 epochs. The right column shows the number of epochs required to reach an accuracy of 0.9 as a function of the number of neurons in a hidden layer for the same learning rates. Each row of plots corresponds to a different bit length (n_bit, ranging from 2 to 5), for the parity check task. The results are averaged over ten independent trials.
Figure 3: Accuracy and convergence characteristics of the EDLA network with the sigmoid AF across various network depths (1, 2, 4, and 8 hidden layers) at a fixed learning rate of 1.0. The left column shows the average classification accuracy as a function of the neurons per hidden layer. The accuracies are averaged from 19900 to 20000 epochs. The right column presents the number of epochs required to reach an accuracy of 0.9 as a function of the number of neurons per hidden layer. Each row of plots corresponds to different bit lengths (n_bit, ranging from 2 to 5). The results are averaged over ten independent trials.
Figure 4: Performance of EDLA with the ReLU AF across various learning rates in the parity check task. The left column shows the accuracy of the EDLA network as a function of neurons in a single hidden layer for different learning rates (LR = 0.01, 0.1, 1.0). The accuracies are averaged from 19900 to 20000 epochs. The right column shows the number of epochs required to achieve an accuracy of 0.9 against the number of neurons in the hidden layer for the same learning rates. Each row of plots corresponds to a different bit length (n_bit), ranging from 2 to 5, for the parity check task. Missing data points in the left column indicate that the accuracy is excessively low. Missing data points in the right column indicate that the accuracy does not reach 0.9. The results are averaged over ten independent trials.
Figure 5: Accuracy and convergence characteristics of the EDLA network with ReLU AF across various network depths (1, 2, 4, and 8 hidden layers), using a fixed learning rate of 0.1. The left column shows average classification accuracy as a function of neurons per hidden layer. The accuracies are averaged from 19900 to 20000 epochs. The right column presents the number of epochs required to reach an accuracy of 0.9 as a function of the number of neurons per hidden layer. Each row of plots corresponds to different bit lengths (n_bit), ranging from 2 to 5. The results are averaged over ten independent trials.
4.2 Regression Tasks
The EDLA network is evaluated on regression tasks using three benchmark datasets: Airfoil Self-Noise, Concrete Compressive Strength, and Energy Efficiency. The Airfoil Self-Noise dataset comprises 1503 samples with five features and a target variable representing the sound pressure level. The Concrete Compressive Strength dataset includes 1030 samples with eight features and a target variable denoting the compressive strength of concrete. The Energy Efficiency dataset comprises 768 samples with eight features and two target variables: heating and cooling loads. In this experiment, only the heating load is employed as the target variable, and the cooling load is not considered. The training and test datasets are randomly split into 80% and 20%, respectively, for all datasets.
The EDLA network consists of an input layer, hidden layers, and an output layer. The neurons in the hidden layers use the sigmoid or ReLU AFs. The neurons in the output layer use the identity AF (i.e., no AF). The training parameters are carefully selected for better performance, with a learning rate of 0.001 for the sigmoid AF and 0.0001 for the ReLU AF. Training is conducted for 500 epochs employing a mini-batch size of 64 samples.
Figure 6 illustrates the Mean Absolute Error (MAE) performance of the EDLA network across varying numbers of neurons (8, 16, 32, 64, 128, 256, 512, and 1024) and different layer configurations (1, 2, and 4 layers). The results indicate that EDLA networks using sigmoid AF with one and four layers achieve low MAE across almost all neuron-count configurations for all datasets. However, the EDLA network with the sigmoid AF does not exhibit performance improvements as the number of neurons or network depth increases.
Notably, the EDLA network with a single hidden layer using the ReLU AF not only achieves low MAE but also consistently outperforms the 1-layer sigmoid network across all three datasets. Furthermore, for this 1-layer network, performance slightly improves as the number of neurons increases. Additionally, increasing the network depth does not yield performance improvements but rather leads to performance degradation compared with shallower architectures. This deterioration is likely caused by the instability of the EDLA network when using ReLU, due to the increase in both neuron count and network depth. As detailed in the caption (Figure 6), this instability was particularly severe for the 4-layer ReLU network, resulting in catastrophically large MAE values (e.g., $10^{30}$) or NaN values for several configurations
These observations collectively suggest that, while the EDLA approach is adequately capable of handling regression tasks, augmenting the number of neurons and network depth does not necessarily improve performance. Furthermore, the EDLA network with the ReLU AF tends to exhibit inferior performance with increased depth. Despite this instability in deeper networks, the ReLU AF demonstrated a slight performance advantage over the sigmoid AF in the most stable (1-layer) configurations.
Figure 6: Performance of the EDLA network on regression tasks across varying layer configurations. Mean Absolute Error (MAE) is shown against the number of neurons per hidden layer for networks with 1, 2, and 4 hidden layers. The rows correspond to the Airfoil Self-Noise (top), Concrete Compressive Strength (middle), and Energy Efficiency (bottom) datasets. The left column shows results for networks using the sigmoid AF (learning rate = 0.001), while the right column shows results for ReLU activation (learning rate = 0.0001). Each point represents the average of four independent trials. Missing data points for the 4-layer ReLU network indicate training instability, where the MAE became excessively large or a NaN (not-a-number) value. The omitted data points (number of neurons, MAE) are as follows: Airfoil: (64, $4.39\times 10^{3}$), (128, $2.66\times 10^{14}$), (256, $1.90\times 10^{31}$), (512, NaN), (1024, NaN); Concrete: (64, $3.09\times 10^{3}$), (128, $3.13\times 10^{13}$), (256, $9.84\times 10^{31}$), (512, $5.79\times 10^{19}$), (1024, NaN); Energy: (64, $1.84\times 10^{4}$), (128, $1.07\times 10^{20}$), (256, NaN), (512, NaN), (1024, NaN).
4.3 Image Classification Tasks
To evaluate the performance of the EDLA network in image classification tasks, the experiments are conducted using the Digits dataset (Optical Recognition of Handwritten Digits), MNIST, Fashion-MNIST, and CIFAR-10 datasets. The Digits dataset consists of 5620 grayscale images with a resolution of $8\times 8$ pixels. Because the Digits dataset is not divided into training and test datasets, 20 % of the entire dataset is randomly selected and used as test data. The MNIST comprises 60,000 training and 10,000 testing grayscale images of handwritten digits, each with a resolution of $28\times 28$ pixels. Similarly, the Fashion-MNIST includes 60,000 training and 10,000 testing grayscale images of Fashion articles at a resolution of $28\times 28$ pixels. The CIFAR-10 consists of 50,000 training and 10,000 testing color images with a resolution of $32\times 32$ pixels. The CIFAR-10 dataset is more complex than the other three datasets because it contains color images. In this experiment, the CIFAR-10 dataset is converted to a flattened one-dimensional vector format.
The EDLA network consists of an input layer, hidden layers, and an output layer. The neurons in the hidden layers use either the sigmoid or ReLU activation functions. The output layer employs the sigmoid AF. To assess its performance across various configurations, the number of neurons in each hidden layer is varied systematically. The learning rates when using the sigmoid and ReLU AFs are 1.0 and 0.1, respectively. The network is trained for 500 epochs when using Digits, MNIST, and Fashion-MNIST datasets, and for 1000 epochs when using the CIFAR-10 dataset. A mini-batch size of 128 samples is used.
Figure 7 illustrates the performance of the EDLA network using the sigmoid AF for various numbers of neurons in each hidden layer. Across all datasets, the classification performance generally improved as the number of neurons in the hidden layers increased. Specifically, for the Digits and MNIST datasets, near-perfect accuracy is rapidly achieved, even with a relatively small number of neurons. For the Fashion-MNIST dataset, while high accuracy is attained, a larger number of neurons contributed to incrementally improved performance. For the CIFAR-10 dataset, the performance is lower compared to the other datasets. This lower performance of CIFAR-10 is attributed to the network architecture, which has a simple layer structure and is not well-suited for processing complex color images as in convolutional neural networks. The performance improvement with an increasing number of neurons is more pronounced in the CIFAR-10 dataset. However, the depth of the EDLA network does not improve performance and can even degrade it. In terms of the reach epoch, for all configurations, the EDLA network shows a sufficiently fast reach of 0.2 and 0.8 accuracy when using the CIFAR-10 dataset and the others. While deeper networks will lead to a slower reach, the number of neurons in each hidden layer will lead to a faster reach.
Similarly, Figure 8 shows the performance of the EDLA network using the ReLU AF across varying numbers of neurons in each hidden layer. The results reveal that the EDLA network with the ReLU AF achieves a performance comparable to that of the sigmoid-based network. However, similar to the sigmoid network, increasing network depth generally leads to worse performance. The trend of accuracy versus neuron count is also more complex: while performance on CIFAR-10 generally improves with more neurons, performance on Digits (4-layer) and Fashion-MNIST (2-layer) slightly degrades. The reach epoch for the ReLU AF also decreases with the number of neurons in each hidden layer. The reach epoch for the ReLU AF is larger (slower) than that for the sigmoid AF.
These observations collectively suggest that EDLA networks can effectively handle relatively simple image classification tasks using both sigmoid and ReLU activation functions. However, more complex datasets such as CIFAR-10 present challenges for the simple EDLA architecture employed in this study. Increasing the number of neurons in each hidden layer generally enhances performance, while increasing network depth does not necessarily yield improvements and may even degrade performance. Regarding convergence speed, increasing the number of neurons generally accelerated convergence for both activation functions. Conversely, increasing network depth tended to slow convergence for sigmoid networks. Furthermore, the reach epoch for the ReLU AF was generally larger (slower) than that for the sigmoid AF across most configurations.
Figure 7: Performance of EDLA with the sigmoid AF on image datasets across layer configurations. The plots in the left column show the accuracy of the EDLA network with the sigmoid AF as a function of the number of neurons in each hidden layer for different layer configurations (1, 2, and 4 layers) across four datasets: Digits, MNIST, Fashion-MNIST, and CIFAR-10. The plots in the right column illustrate the number of epochs required to achieve an accuracy of $\theta_{\mathrm{threshold}}$ against the number of neurons in each hidden layer for the same layer configurations. For CIFAR-10, $\theta_{\mathrm{threshold}}$ is set to 0.2. Otherwise, $\theta_{\mathrm{threshold}}$ is set to 0.8. The results are averaged over four independent trials.
Figure 8: Performance of EDLA with the ReLU AF on image datasets across layer configurations. The plots in the left column show the accuracy of the EDLA network with the ReLU AF as a function of the number of neurons in each hidden layer for different layer configurations (1, 2, and 4 layers) across four datasets: Digits, MNIST, Fashion-MNIST, and CIFAR-10. The plots in the right column illustrate the number of epochs required to achieve an accuracy of $\theta_{\mathrm{threshold}}$ against the number of neurons in each hidden layer for the same layer configurations. For CIFAR-10, $\theta_{\mathrm{threshold}}$ is set to 0.2. Otherwise, $\theta_{\mathrm{threshold}}$ is set to 0.8. The results are averaged over four independent trials.
4.4 Analysis of Internal Representations
In this subsection, we investigate the internal representations learned by the EDLA network by analyzing the synaptic weights in the hidden layers. The experiments use an EDLA network with a single hidden layer containing two neurons with sigmoid AFs. The network is trained on the Digits dataset over 500 epochs.
Figure 9 shows a heatmap of the synaptic weights connecting the input layer to the hidden layer within the EDLA network, specifically trained to classify the digit “zero” from handwritten digits. The heatmap reveals that the network learned meaningful features from the input data. Notably, the neurons in the hidden layer exhibit a strong response to specific patterns within the images. For instance, the weights connecting the positive neurons in the input layer to the positive neurons in the hidden layer (denoted as "P-P" connections) show high activation, shaping the digit “zero.” This result suggests that the EDLA method effectively enables neurons to capture and represent the relevant digit-specific features. Additionally, similar “zero-shaped” patterns also emerge in other synaptic connections.
Synaptic weight distributions in neural networks provide valuable insights into the learning behavior of the network and features extracted from the input data. To explore this further, we analyze the distribution of weights across the EDLA network. Figure 10 shows a histogram representing the synaptic weights within the trained EDLA network. The distribution appears to be predominantly centered around small magnitudes, indicating that most weights are relatively small. However, a distinct spread is observed with a smaller subset of synapses exhibiting large positive or negative values. This broad distribution suggests that the EDLA update unilaterally increases the absolute values of the weights while balancing the excitatory and inhibitory weights.
Figure 9: A heatmap of the synaptic weights connecting the input layer to the first hidden layer in the EDLA network, trained to classify the digit “zero” in the Digits dataset. Each subplot shows the weight patterns for individual connection patterns, labeled according to connection types: P-P (positive input to positive hidden sublayers), N-P (negative input to positive hidden sublayers), P-N (positive input to negative hidden sublayers), and N-N (negative input to negative hidden sublayers). Color gradients denote weight values, with yellow indicating strongly positive weights and purple indicating strongly negative weights.
Figure 10: Histogram of synaptic weight distribution in the EDLA Network. The horizontal axis denotes the synaptic weight values, and the vertical axis indicates the corresponding frequencies.
4.5 Stabilization of Learning Dynamics in EDLA with ReLU Activation Function
Deeper EDLA networks employing the ReLU AF often exhibit severe performance degradation when the number of hidden layers (depth) increases. This degradation arises from the unbounded nature of the ReLU AF, which can generate excessively large activations and thereby produce excessively large and unstable weight updates during training. To mitigate these effects, two stabilization techniques are introduced: the use of small-scale initial weights and RMS normalization of neuronal activations (Sec. 2.7). The former limits the magnitude of initial activations, while the latter normalizes neuronal pre-activations based on their root-mean-square (RMS) values. These mechanisms jointly stabilize the learning dynamics by suppressing excessive activations and preventing abrupt weight updates.
Figure 11 presents the results of EDLA networks with the ReLU AF under these stabilization strategies across a range of network widths and depths for both regression and image classification tasks. For regression (Airfoil and Concrete datasets) and classification (MNIST and CIFAR-10 datasets), the learning rates are set to 0.0001 and 0.1, respectively. The weights are drawn from uniform distributions over $0,\mathrm{IS}]$ and $\mathrm{IS},0]$ for synapses connecting same-type and different-type neurons, respectively. The initial scale (IS) was 0.0001 for the small-scale variant, and 1.0 for both the standard and RMS normalization. RMS normalization is applied to each layer except the input and output layers. The mini-batch sizes are 64 for regression and 128 for classification.
For regression tasks, both small initial weights and RMS normalization suppress extreme or NaN (not-a-number) MAE values that arise in deeper networks, leading to stable training. For classification tasks, small initial weights do not significantly influence the accuracy of shallow networks (1–2 layers), but in deeper configurations (4 layers) on MNIST, they tend to degrade accuracy slightly. In contrast, RMS normalization improves accuracy in deeper architectures, particularly for the CIFAR-10 dataset with larger hidden dimensions. These findings indicate that small initial weights and RMS normalization are effective in preventing instability in deeper EDLA networks with the ReLU AF, stabilizing training without compromising general performance.
Figure 11: Performance comparison of EDLA with ReLU AF using two stabilization techniques: small initial weights (IS = 0.0001, left column) and RMS normalization (right column). The plots show performance across varying layer configurations (1, 2, and 4 layers) and neuron counts. Top two rows (Regression): MAE on Airfoil and Concrete datasets (LR = 0.0001). Bottom two rows (Classification): Accuracy on MNIST and CIFAR-10 datasets (LR = 0.1). Results are averaged over four independent trials.
The observed improvement in deeper EDLA networks likely arises from the suppression of exploding activations by these stabilization techniques. The instability of ReLU-based EDLA networks can primarily be attributed to two factors: (i) exploding activations, and (ii) the so-called dying ReLU phenomenon, in which neurons become inactive and cease to contribute to learning. The former refers to the uncontrolled growth of activations during forward propagation, which induces large Hebbian-like weight updates under EDLA’s learning rule. This effect becomes increasingly pronounced as the depth of the network grows, resembling the well-known exploding gradient phenomenon in recurrent or very deep networks [31, 32]. The latter, known as the dying ReLU problem [33, 34], occurs when neurons become inactive (outputting zero for all inputs), preventing them from contributing to learning. In practice, these two effects interact: large activations in early training produce extreme weight updates, which then push many neurons into permanently inactive states, further deteriorating the network’s representational capacity.
To elucidate the effects of these stabilization methods and clarify the cause of instability, several internal metrics were monitored. Test MAE was evaluated on the test set to track generalization performance. Concurrently, several internal dynamics were monitored on the training data, including the network-wide maximum and mean activations, per-layer activation maxima, the network-wide maximum and mean absolute weight updates, and the global dead-unit ratio (the percentage of inactive neurons). In this study, a neuron was considered dead if its activation remained below the fixed threshold of 0.01 for more than 95% of training samples. Figure 12 compares these metrics across three EDLA variants: the baseline EDLA with the ReLU AF, the EDLA with small-scale initialization (IS = 0.0001), and the EDLA with RMS normalization. The initial scale (IS) was set to 1.0 for both the baseline EDLA and the EDLA with RMS normalization. This internal analysis was conducted on the Concrete regression task using a network with four hidden layers of 32 neurons each, a learning rate of 0.0001, and a mini-batch size of 64.
As shown in Figure 12A, the baseline EDLA exhibits large and unstable test MAE values throughout training, indicating poor convergence and sensitivity to parameter updates. In contrast, both small initial weights and RMS normalization suppress large fluctuations in MAE, leading to smoother and smaller test errors. The small-scale initialization achieves the lowest MAE overall, while RMS normalization provides similarly stable and improved performance. Figures 12B and 12C show that the baseline network experiences a rapid increase in the maximum activation magnitude early in training. The small-scale initialization moderates this activation growth, and RMS normalization stabilizes activations at moderate levels throughout training. Interestingly, the RMS-normalized network maintains higher mean activations than the other variants, even when its maximum activations are well controlled, suggesting that RMS normalization prevents neuron inactivity while maintaining balanced activation levels. This interpretation is supported by the lower dead-unit ratio observed in Figure 12I.
The per-layer activation trends in Figures 12D–F reveal that deeper layers tend to exhibit larger activations due to the cumulative effect of the unbounded ReLU activation, while both stabilization techniques attenuate this amplification across all layers. Figures 12G and 12H further show that the baseline EDLA undergoes excessively large weight updates in the early stages of training. In the standard EDLA, although the mean weight update is smaller than that of the stabilized variants, the weights of many neurons remain nearly unchanged, while a few experience extremely large updates, leading to unstable learning dynamics. Consequently, the global dead-unit fraction shown in Figure 12I reveals that the baseline model suffers from a high proportion of inactive neurons. Many neurons become permanently dead early in training, severely limiting the network’s representational capacity. In contrast, both stabilization methods maintain smaller and more consistent weight update magnitudes across epochs. Both stabilization methods substantially reduce the occurrence of dead units, with RMS normalization showing the most pronounced effect. These results indicate that the stabilization techniques effectively regulate activation and weight dynamics.
These results collectively demonstrate that the instability of ReLU-based EDLA networks arises from the combined effects of exploding activations and dead ReLU units. Small initial weights alleviate early activation explosions, whereas RMS normalization further stabilizes learning by controlling both activation magnitudes and update scales. Importantly, RMS normalization maintains sufficient neuron activity and prevents the collapse of representational diversity. Overall, these findings indicate that appropriate control of activation magnitudes is essential for stable training in EDLA networks. The effectiveness of RMS normalization in particular parallels the success of normalization techniques such as Batch Normalization [26] and Layer Normalization [25], highlighting that regulation of internal activation dynamics is a key factor for achieving stable and efficient learning in non-standard architectures such as EDLA.
Figure 12: Analysis of internal dynamics for three EDLA-ReLU variants on the Concrete regression task ($n_{\mathrm{hidden}}=32$, $L=4$, LR = 0.0001). Variants are: ReLU (scale = 1.0) (baseline EDLA), ReLU (scale = 0.0001) (EDLA with the small initial scale), and RMS Norm + ReLU (scale = 1.0) (EDLA with RMS normalization). Metrics shown: (A) Test MAE; (B) network-wide maximum activation; (C) network-wide mean activation; (D–F) per-layer post-activation maxima (indices 1–4 correspond to hidden layers); (G) network-wide maximum absolute weight update; (H) network-wide mean absolute weight update; (I) global dead-unit fraction. Metric (A) was evaluated on the test set, while metrics (B)–(I) were evaluated on the training set. (D–H) are plotted on a log10 scale ($\varepsilon=10^{-12}$ added to avoid $\log(0)$).
4.6 Comparison with multilayer perceptron trained with backpropagation
To benchmark the performance of the EDLA network, it is compared with a standard multilayer perceptron (MLP) trained using backpropagation. The MLP architecture mirrors that of the EDLA network, consisting of an input layer, hidden layers, and an output layer. The neurons in the hidden layers utilize either the sigmoid or ReLU AFs, while the output layer employs the identity AF for regression tasks and the sigmoid AF for classification tasks. The MLP is trained using the AdamW optimizer (without weight decay), with default parameters except for the learning rate. The learning rates are set to 0.001 for the MLP with the sigmoid AF and 0.0001 for the MLP with the ReLU AF in regression tasks. For classification tasks, the learning rates are set to 0.01 for the MLP with the sigmoid AF and 0.001 for the MLP with the ReLU AF. Training is conducted for 500 epochs with a mini-batch size of 64 for regression tasks and 128 for classification tasks. For CIFAR-10, training is conducted for 1000 epochs with a mini-batch size of 128. The training configurations of EDLA variants are the same as in previous sections.
Table 1 presents the performance comparison across four representative datasets. In the “Model (Act)” column, the activation function and stabilization variant are indicated in parentheses. “IS” refers to the initial scale parameter used for weight initialization, and “RMS” indicates the application of RMS normalization to the pre-activations. For example, “EDLA (ReLU, IS=0.0001)” represents the ReLU-based EDLA with a small initial scale of 0.0001, and “EDLA (ReLU+RMS)” represents the ReLU-based EDLA incorporating RMS normalization.
For a fair comparison, the parameter counts of EDLA and MLP in each configuration were matched or approximated as closely as possible. The results confirm that the backpropagation-trained MLP is a more powerful optimizer, consistently achieving superior performance. However, the magnitude of this performance gap is highly dependent on task complexity and network depth. On simpler, shallower tasks, EDLA is highly competitive. On the 1-layer MNIST task, for instance, the best EDLA variant (Sigmoid) achieves 0.973 accuracy, only slightly trailing the MLP’s 0.981. In contrast, the gap widens significantly on deeper, more complex tasks. On the 4-layer Airfoil regression task, the MLP (ReLU) achieves an MAE of 1.599, whereas the best-performing stabilized EDLA (ReLU, Small IS) only reaches 6.381. This highlights a key finding: while the stabilization techniques (Sec. 4.5) successfully prevented the catastrophic failure of deep EDLA-ReLU networks (which produced MAE $10^{30}$), they were not sufficient to close the fundamental optimization gap with a standard backpropagation-trained MLP.
This performance gap, however, is expected and illustrates the fundamental trade-offs of the EDLA framework. Backpropagation achieves its high performance by propagating exact, layer-specific error gradients, allowing for fine-grained optimization across the entire network. EDLA, by design, sacrifices this optimization power for biological plausibility, relying instead on an approximated, single global error signal diffused uniformly to all layers. Furthermore, this gap is compounded on multi-class tasks (MNIST, CIFAR-10) by EDLA’s multi-head architecture. An MLP uses a shared representational body and a single final layer for all classes. In contrast, the EDLA implementation must train $K$ independent networks, one for each output class. This approach not only requires significantly more parameters (as analyzed in Sec. 2.6) but also prevents the network from learning shared, general-purpose features across classes, likely degrading performance on high-dimensional, multi-class problems.
On the CIFAR-10 dataset, the MLP’s performance is also relatively low, but this result is not irregular for a simple MLP architecture. Simple MLPs with ReLU show about 50% accuracy on CIFAR-10. For example, a simple MLP with ReLU activation, six layers, and 1024 neurons per layer achieves 54.2% accuracy with Layer Normalization on CIFAR-10 [35], while a two-layer MLP (width 512, ReLU, and 0.2 dropout) achieves 54–58% accuracy [36]. Thus, the MLP’s performance in the experiments is consistent with the findings in the literature, confirming that both EDLA and traditional MLPs encounter similar architectural limitations when applied to high-dimensional image data without convolutional priors.
Table 1: Performance comparison (MNIST, CIFAR-10, Airfoil, Concrete). For classification tasks the metric is accuracy; for regression tasks the metric is mean absolute error (MAE). The best performance for each dataset and architecture class (e.g., 4-layer) is highlighted in bold. Note that the parameter counts for EDLA and MLP include biases.
Dataset Metric Model (Act) Arch (L x N) Params Value (Mean)
Airfoil MAE EDLA (Sigmoid) 1 x 256 7.2k 3.828
MAE EDLA (ReLU) 1 x 256 7.2k 3.439
MAE EDLA (ReLU, IS=0.0001) 1 x 256 7.2k 3.751
MAE EDLA (ReLU+RMS) 1 x 256 7.2k 3.259
MAE MLP (Sigmoid) 1 x 1024 7.2k 2.411
MAE MLP (ReLU) 1 x 1024 7.2k 2.739
MAE EDLA (Sigmoid) 4 x 256 1.58M 9.072
MAE EDLA (ReLU) 4 x 256 1.58M $1.90\times 10^{31}$
MAE EDLA (ReLU, IS=0.0001) 4 x 256 1.58M 6.381
MAE EDLA (ReLU+RMS) 4 x 256 1.58M 20.941
MAE MLP (Sigmoid) 4 x 768 1.78M 5.842
MAE MLP (ReLU) 4 x 768 1.78M 1.599
Concrete MAE EDLA (Sigmoid) 1 x 256 10.2k 7.832
MAE EDLA (ReLU) 1 x 256 10.2k 7.511
MAE EDLA (ReLU, IS=0.0001) 1 x 256 10.2k 8.357
MAE EDLA (ReLU+RMS) 1 x 256 10.2k 6.271
MAE MLP (Sigmoid) 1 x 1024 10.2k 4.349
MAE MLP (ReLU) 1 x 1024 10.2k 4.630
MAE EDLA (Sigmoid) 4 x 256 1.59M 7.986
MAE EDLA (ReLU) 4 x 256 1.59M $9.84\times 10^{31}$
MAE EDLA (ReLU, IS=0.0001) 4 x 256 1.59M 19.814
MAE EDLA (ReLU+RMS) 4 x 256 1.59M 14.236
MAE MLP (Sigmoid) 4 x 768 1.78M 13.722
MAE MLP (ReLU) 4 x 768 1.78M 3.084
MNIST Accuracy EDLA (Sigmoid) 1 x 16 503.1k 0.973
Accuracy EDLA (ReLU) 1 x 16 503.1k 0.969
Accuracy EDLA (ReLU, IS=0.0001) 1 x 16 503.1k 0.970
Accuracy EDLA (ReLU+RMS) 1 x 16 503.1k 0.971
Accuracy MLP (Sigmoid) 1 x 512 407.1k 0.981
Accuracy MLP (ReLU) 1 x 512 407.1k 0.981
Accuracy EDLA (Sigmoid) 4 x 64 3.00M 0.968
Accuracy EDLA (ReLU) 4 x 64 3.00M 0.949
Accuracy EDLA (ReLU, IS=0.0001) 4 x 64 3.00M 0.942
Accuracy EDLA (ReLU+RMS) 4 x 64 3.00M 0.962
Accuracy MLP (Sigmoid) 4 x 1024 3.96M 0.980
Accuracy MLP (ReLU) 4 x 1024 3.96M 0.980
CIFAR-10 Accuracy EDLA (Sigmoid) 1 x 32 3.93M 0.421
Accuracy EDLA (ReLU) 1 x 32 3.93M 0.410
Accuracy EDLA (ReLU, IS=0.0001) 1 x 32 3.93M 0.364
Accuracy EDLA (ReLU+RMS) 1 x 32 3.93M 0.387
Accuracy MLP (Sigmoid) 1 x 1024 3.16M 0.471
Accuracy MLP (ReLU) 1 x 1024 3.16M 0.536
Accuracy EDLA (Sigmoid) 4 x 16 2.03M 0.193
Accuracy EDLA (ReLU) 4 x 16 2.03M 0.216
Accuracy EDLA (ReLU, IS=0.0001) 4 x 16 2.03M 0.243
Accuracy EDLA (ReLU+RMS) 4 x 16 2.03M 0.353
Accuracy MLP (Sigmoid) 4 x 512 2.37M 0.451
Accuracy MLP (ReLU) 4 x 512 2.37M 0.541
5 Conclusion and Discussion
The Error Diffusion Learning Algorithm (EDLA) provides a biologically inspired alternative to conventional backpropagation methods for training neural networks and is characterized by the diffusion of a global error signal across network layers. This mechanism is analogous to reward-based learning processes observed in biological neural systems. Moreover, EDLA mirrors neuromodulatory systems, such as the dopaminergic system, which projects broadly across cortical and subcortical areas, disseminating their signals globally [12]. In contrast to traditional MLPs, EDLA incorporates structured interactions between neurons with excitatory and inhibitory synapses. This reflects the biologically observed balance between these types of synapses.
However, EDLA is inherently constrained by its single-output architecture, limiting its applicability to multiclass classification tasks. A practical solution involves parallelizing multiple EDLA networks, although this approach increases computational costs. This parallelized architecture mirrors the modular structure of biological neural systems in which specialized circuits process distinct features or tasks independently. In such a configuration, each EDLA network separately optimizes its own output according to a unique global error signal to that network. This modular structure aligns with distributed processing mechanisms observed in biological neural networks.
EDLA demonstrates efficacy across various tasks, including parity-check, regression, and image classification. Experimental evaluations indicate that the EDLA achieves satisfactory performance across various datasets. Notably, the ReLU AF achieves a performance comparable to that of the sigmoid function originally proposed by Kaneko [9], where the ReLU often converges more quickly. The faster convergence of ReLU networks observed here replicates earlier findings showing that the ReLU AF accelerates learning in neural networks [37].
In contrast, the experiments in this study consistently showed that increasing EDLA’s network depth, unlike standard deep learning, does not reliably improve performance and can even lead to degradation (Figs. 6, 7, 8). We hypothesize this stems from EDLA’s unique update mechanism based on a single global error signal. Unlike backpropagation, where gradients can vanish in deep layers due to the chain product of activation function derivatives [38, 32], EDLA’s global signal ensures all weights receive an update signal. While this avoids the classical vanishing gradient problem caused by the chain product of derivatives, it does not guarantee effective credit assignment for deep hierarchies. Updates driven by the single scalar error $d$ might not be optimal for each specific layer’s needs. Furthermore, changes induced in lower-layer activations propagate upwards and can be amplified through successive layers. This amplification becomes particularly severe (“exploding activations”) when using unbounded activation functions like ReLU. However, even with bounded functions like sigmoid, the cumulative effect of layer-by-layer amplification driven by a non-specific global signal might interfere with learning the complex, hierarchical features typically expected from deeper networks, potentially explaining the lack of performance improvement or degradation. This amplification effect is inherently less pronounced in shallower networks. Consequently, for EDLA, increasing network width appears to be a more effective and stable strategy for improving performance than increasing depth.
A key finding was the depth-related instability observed, especially with the ReLU activation function. This instability manifested severely in deep ReLU regression networks, causing catastrophic failures (e.g., NaN MAE). This failure mode is likely exacerbated by the regression task’s linear output layer: exploding activations from the final hidden layer directly produce an extremely large output $y$, which in turn generates an enormous global error signal d=tyd=t-y. This huge error signal then catastrophically amplifies weight updates throughout the entire network. This depth-related instability primarily stems from the unbounded nature of the ReLU AF itself, a phenomenon observed in standard deep networks as well [31, 32]. As signals propagate through multiple layers, ReLU’s lack of saturation allows activation magnitudes to grow excessively (“exploding activations”). While this is a general challenge in deep ReLU networks, it becomes particularly problematic within the EDLA framework. EDLA’s Hebbian-like learning rule translates these large activations directly into excessively large weight updates. The single global error signal, while perhaps not the primary initiator of the explosion, likely contributes to propagating instability across the network once large activations emerge.
Crucially, the internal dynamics analysis (Figure 12) provides direct quantitative evidence for this cascade in EDLA, but reveals a mechanism for vanishing updates different from classical vanishing gradient. The baseline deep ReLU network exhibited exploding maximum activations (Fig. 12B) and extreme initial weight updates (Fig. 12G, H), immediately followed by a dramatic rise in the dead neuron fraction to over 80% (Fig. 12I). This strongly indicates that performance collapse stems from ReLU-driven exploding activations, exacerbated by EDLA’s update rule, triggering widespread neuron death, which is itself a primary source of vanishing weight updates, as dead units provide zero update signals. In contrast, the saturating sigmoid function inherently limits activation magnitudes and thus update sizes, explaining its greater stability in EDLA.
It was demonstrated that this instability can be effectively managed. While smaller learning rates and initial weights help, RMS normalization proved particularly robust, controlling extreme activations while crucially maintaining neuron activity (low dead unit fraction, Fig. 12I), thus enabling stable learning in deeper ReLU networks.
Despite successful stabilization, comparison with backpropagation-trained MLPs (Table 1) revealed a persistent performance gap, widening with network depth and task complexity. Although EDLA achieved highly competitive results in shallow configurations, MLPs consistently performed better overall. This underscores EDLA’s fundamental trade-off: sacrificing the optimization power of exact, layer-specific gradients for the biological plausibility of an approximated, global error signal. The multi-head architecture likely exacerbates this gap in multi-class settings by preventing shared feature learning.
Nonetheless, EDLA remains a significant contribution towards biologically realistic learning algorithms. Its reliance on local activity and a global signal holds promise for neuromorphic computing and insights into brain function.
Future work could explore several avenues to enhance EDLA’s capabilities. First, integrating eligibility traces could enable single EDLA networks to handle multiclass classification tasks, reducing the need for parallel architectures. Eligibility traces could allow neurons to maintain a memory of past activations, facilitating more complex decision boundaries. However, determining which neurons should maintain these traces remains an open question. Second, investigating hybrid architectures that combine EDLA with backpropagation in a complementary manner might leverage the strengths of both approaches. For instance, EDLA might be employed for initial coarse learning phases, followed by fine-tuning via backpropagation. Finally, applying EDLA to other architectures such as recurrent and convolutional neural networks [39, 40, 41] could extend its applicability to temporal and spatial data domains, respectively.
Acknowledgments
The author gratefully acknowledges the assistance of Chat AI assistants (OpenAI ChatGPT and Google Gemini) and grammar checking service (Grammarly) in improving the grammar and style of this work.
Data Availability
The source code generated during this study is publicly available on GitHub at https://github.com/KazuhisaFujita/EDLA. The datasets analysed during this study are publicly available in the following repositories:
Funding Declaration
The author received no funding for this work.
Ethics declarations
Conflict of Interest
The author declares no competing interests.
Ethical Approval
This article does not contain any studies with human participants or animals performed by the author.
Appendix (for pedagogical completeness)
Appendix A Derivation of Backpropagation in a Multilayer Perceptron
A multilayer perceptron (MLP) is a class of feedforward artificial neural networks characterized by multiple interconnected neuron layers: an input layer, one or more hidden layers, and an output layer. Typically, each neuron within one layer is fully connected to all neurons in the subsequent layer.
The input to an MLP is represented by the vector $\mathbf{x}=[x_{0},x_{1},\ldots,x_{N}]^{\mathrm{T}}\in\mathbb{R}^{N+1}$, where $x_{0}$ commonly serves as a bias term (usually set as $x_{0}=1$) and $N$ denotes the number of input features. This input vector is received by the input layer and subsequently propagated forward to the first hidden layer.
For neuron $j$ in the first hidden layer $(l=1)$, the activation $a_{j}^{(1)}$ is calculated as the weighted sum of its inputs:
$a_{j}^{(1)}=\mathbf{w}_{j}^{(1)\mathrm{T}}\mathbf{x}$ (28)
where $\mathbf{w}_{j}^{(1)}=[w_{j0}^{(1)},w_{j1}^{(1)},\ldots,w_{jN}^{(1)}]^{\mathrm{T}}$ represents the weight vector connecting the input layer to neuron $j$ in the first hidden layer $(l=1)$. Here, $w_{ji}^{(1)}$ denotes the weight connecting neuron $i$ in the input layer to neuron $j$ in the first hidden layer. The output of neuron $j$ is then obtained by applying the activation function $g(\cdot)$:
$z_{j}^{(1)}=g(a_{j}^{(1)}).$ (29)
For subsequent hidden layers (l=2,,L)(l=2,...,L), the activation $a_{j}^{(l)}$ for neuron $j$ in layer $l$ is computed as follows:
$a_{j}^{(l)}=\mathbf{w}_{j}^{(l)\mathrm{T}}\mathbf{z}^{(l-1)},$ (30)
where $\mathbf{z}^{(l-1)}=[z_{0}^{(l-1)},z_{1}^{(l-1)},\ldots,z_{M^{(l-1)}}^{(l-1)}]^{\mathrm{T}}\in\mathbb{R}^{M^{(l-1)}+1}$ is the output vector of layer l1l-1, $M^{(l-1)}$ is the number of neurons in this layer, and $z_{0}^{(l-1)}$ corresponds to the constant output of a bias neuron fixed to one. $w_{ji}^{(l)}$ denotes the weight connecting neuron $i$ in layer l1l-1 to neuron $j$ in layer $(l)$. The output for neuron $j$ is then given by
$z_{j}^{(l)}=g(a_{j}^{(l)}).$ (31)
Finally, the output for neuron $k$ in the output layer $(L+1)$ is similarly computed as:
$y_{k}=f(a_{k}^{(L+1)})=f(\mathbf{w}_{k}^{(L+1)\mathrm{T}}\mathbf{z}^{(L)})$ (32)
where $f(\cdot)$ is the activation function for the output neurons, and $\mathbf{w}_{k}^{(L+1)}$ is the weight connecting the neurons in the final hidden layer $(L)$ to output neuron $k$.
Training an MLP seeks the optimal weights $\mathbf{w}$ that minimize a predetermined loss function $E(\mathbf{w})$, such as mean squared error or cross-entropy loss. Gradient descent is a widely used optimization method, in which weights are iteratively updated in the direction that locally reduces the value of $E(\mathbf{w})$. To efficiently compute gradients necessary for this optimization, the backpropagation algorithm employs the chain rule, propagating the error signals sequentially from the output layer backward through each hidden layer to the input layer.
Consider an output neuron $k$ in the output layer $(L+1)$. Using the chain rule, the partial derivative of the loss function $E(\mathbf{w})$ with respect to the weight $w_{kj}^{(L+1)}$ connecting neuron $j$ in the last hidden layer $(L)$ to neuron $k$ in the output layer is computed as follows:
$\frac{\partial E(\mathbf{w})}{\partial w_{kj}^{(L+1)}}=\frac{\partial E(\mathbf{w})}{\partial a_{k}^{(L+1)}}\frac{\partial a_{k}^{(L+1)}}{\partial w_{kj}^{(L+1)}}=\delta_{k}^{(L+1)}z_{j}^{(L)}$ (33)
where
$\delta_{k}^{(L+1)}\equiv\frac{\partial E(\mathbf{w})}{\partial a_{k}^{(L+1)}}=f^{\prime}(a_{k}^{(L+1)})\frac{\partial E(\mathbf{w})}{\partial y_{k}}.$ (34)
This quantity, often referred to as the error for output neuron $k$, is the partial derivative of the loss $E(\mathbf{w})$ with respect to the neuron’s activation $a_{k}^{(L+1)}$. By the chain rule, this derivative can be factored into two components: the partial derivative of the loss function with respect to the neuron’s output, $\frac{\partial E(\mathbf{w})}{\partial y_{k}}$, and the derivative of the activation function evaluated at $a_{k}^{(L+1)}$, namely $f^{\prime}(a_{k}^{(L+1)})$.
To calculate the error $\delta_{j}^{(l)}$ for hidden layer $(l)$, we make use of the chain rule again. The partial derivative of the loss function with respect to the weight $w_{ji}^{(l)}$ is:
$\frac{\partial E(\mathbf{w})}{\partial w_{ji}^{(l)}}=\frac{\partial E(\mathbf{w})}{\partial a_{j}^{(l)}}\frac{\partial a_{j}^{(l)}}{\partial w_{ji}^{(l)}}=\delta_{j}^{(l)}z_{i}^{(l-1)}.$ (35)
The error $\delta_{j}^{(l)}$ is defined as:
$\delta_{j}^{(l)}\equiv\frac{\partial E(\mathbf{w})}{\partial a_{j}^{(l)}}=g^{\prime}(a_{j}^{(l)})\sum_{k}\delta_{k}^{(l+1)}w_{kj}^{(l+1)},$ (36)
where $g^{\prime}(\cdot)$ represents the derivative of the activation function $g(\cdot)$ evaluated at $a_{j}^{(l)}$. Intuitively, the error at hidden neuron $j$ is computed based on the weighted sum of error signals from the neurons in the subsequent layer $(l+1)$, thus sequentially propagating the error signal backward through the network.
Once the partial derivatives $\frac{\partial E(\mathbf{w})}{\partial w_{ji}^{(l)}}$ are computed for all weights, the update rule in standard gradient descent is:
$w_{ji}^{(l)}\leftarrow w_{ji}^{(l)}-\eta\frac{\partial E(\mathbf{w})}{\partial w_{ji}^{(l)}},$ (37)
where $\eta$ is the learning rate. This iterative cycle—consisting of a forward pass (computation of neuron activations), a backward pass (error computation), and weight updates—continues until a predetermined convergence criterion is satisfied or a specified number of training epochs has been reached.
The gradient of the loss function with respect to the weights is computed using training data. In practice, to compute the gradient, most practitioners use a method called stochastic gradient descent (SGD) [10]. In SGD, the calculation of the gradient is performed on a subset of the training data, known as a mini-batch. This approach enables efficient computation of the gradient, making it feasible to train large neural networks using extensive datasets.
Traditional MLPs have an advantage when handling network architectures requiring multiple outputs. MLPs can simultaneously optimize the weights associated with each output. This is attributed to the backpropagation mechanism in MLPs, where the weights connecting the output layer to the preceding hidden layer are specifically adjusted to minimize the error associated with each output. Concurrently, hidden-layer neurons are trained to extract and generalize pertinent features from the input data. This hierarchical optimization enables MLPs to efficiently model complex input-output relationships and successfully generalize the learned features.